### v1

# Robust Few-shot Transfer Learning for Knowledge Base Question Answering with Unanswerable Questions

Abstract
Real-world KBQA applications require models that are (1) robust – e.g., can differentiate between answerable and unanswerable questions, and (2) low-resource – do not require large training data. Towards this goal, we propose the novel task of few-shot transfer for KBQA with unanswerable questions. We present FUn-FuSIC that extends the state-of-the-art (SoTA) few-shot transfer model for answerable-only KBQA to handle unanswerability. It iteratively prompts an LLM to generate logical forms for the question by providing feedback using a diverse suite of syntactic, semantic and execution guided checks, and adapts self-consistency to assess confidence of the LLM to decide answerability. Experiments over newly constructed datasets show that FUn-FuSIC outperforms suitable adaptations of the SoTA model for KBQA with unanswerability, and the SoTA model for answerable-only few-shot-transfer KBQA.
1 Introduction
The semantic parsing formulation of the Knowledge Base Question Answering (KBQA) task takes a Knowledge Base (KB) and a natural language question as input and requires as output a logical form (or program) that produces an answer upon execution over the KB. KBQA has important real-world applications, which require KBQA systems to be low-resource (requiring only a few task-specific labeled examples) and robust. One important aspect of robustness is the ability to identify unanswerable questions, which arise naturally in real applications Patidar et al. (2023). Traditional supervised models Saxena et al. (2022); Zhang et al. (2022); Mitra et al. (2022); Wang et al. (2022); Das et al. (2022); Ye et al. (2022); Chen et al. (2021); Das et al. (2021); Shu et al. (2022); Gu et al. (2023) and even recent LLM-based few-shot in-context learning (FS-ICL) architectures Li et al. (2023); Nie et al. (2024) are found wanting in both aspects.
Recent work has addressed low-resource KBQA in the form of few-shot transfer learning Patidar et al. (2024) – the idea is to exploit existing KBQA datasets (source) and transfer them few-shot to a target setting. Their FuSIC-KBQA architecture uses supervised KB retrieval followed by generation and refinement of logical forms using LLM few-shot in-context learning.
A parallel direction addresses unanswerable questions for KBQA Patidar et al. (2023). Unanswerability arises due to deficiencies in the schema or the data in the KB.  Faldu et al. (2024) proposes RetinaQA – a supervised architecture (with high training data needs) that is robust to both answerable and unanswerable questions. However, no existing architecture simultaneously addresses the low-resource and robustness requirements.
In response, we propose the task of few-shot transfer learning for KBQA in the unaswerability setting. Specifically, the target KBQA task offers only a few labeled examples, while the source KBQA task has thousands of labeled examples containing answerable questions only.
Building a model for this setting could involve extending FuSIC-KBQA for unanswerability or RetinaQA for few-shot transfer. Since RetinaQA’s design requires large volumes of labeled training data in the target and it fares poorly in few-shot setting, we build our model on top of FuSIC-KBQA. A natural vanilla extension is to modify the LLM prompt for unanswerability and including different types of unanswerable questions in the few shots, as proposed in the broader unanswerable QA literature Slobodkin et al. (2023). Unfortunately, we find that this largely fails for unanswerable questions that do not have a valid logical form – the simple modification regularly labels these as answerable and outputs incorrect logical forms. A second issue is that FuSIC-KBQA prompts LLM to refine the logical form if its execution leads to an empty answer – but this assumes that all questions are answerable and naturally breaks down in the face of unanswerability.
To handle this, we develop FUn-FuSIC (Feedback for Unanswerability in FuSIC-KBQA). FUn-FuSIC takes the vanilla approach further by detecting when the LLM is very uncertain about the logical form, which is suggestive of the question being unanswerable. For this, it employs logical form refinement by providing LLM repeated feedback about potential errors, including syntactic, semantic and execution-related error checks, and a novel strategy of back-translating the logical form to check equivalence with the original question. To assess the model’s lack of certainty about the logical form, FUn-FuSIC adapts self-consistency Wang et al. (2023) – in the absence of enough certainty, schema-level unanswerability is inferred. On the other hand, when a sufficient number of logical forms agree on the answer (empty or non-empty), the LLM is finally prompted to choose the best one, along with inferring answerability (for non-empty answer) or data-related unanswerability (for empty answer).
For evaluation, we first create two new datasets for few-shot transfer for KBQA with unanswerability. Using these, we show that, in aggregate, FUn-FuSIC comprehensively outperforms RetinaQA, when adapted for few-shot transfer, as well as the vanilla extension to FuSIC-KBQA.
We further demonstrate that the broader idea of iterative generation of logical forms, with feedback from a diverse suite of error checks, holds promise for KBQA in general. Using experiments over existing few-shot transfer datasets for answerable KBQA, we find that the restriction of FUn-FuSIC for the answerable setting outperforms FuSIC-KBQA, which is the SoTA KBQA model in this setting.
In summary, our specific contributions are as follows. (a) We propose the problem of few-shot transfer for KBQA with unanswerability. (b) We present FUn-FuSIC that uses iterative feedback for error checks and self-consistency adapted for unaswerability. (c) We create new datasets for the proposed task which we make public  [^1] https://anonymous.4open.science/r/FUn-FuSIC-1704. (d) We show that FUn-FuSIC outperforms adaptation of SoTA KBQA models for our setting. (e) We also show that in the standard answerable setting, FUn-FuSIC outperforms the corresponding SoTA model.
2 Related Work
There is a lot of work on in-domain KBQA using supervised models  Ye et al. (2022); Chen et al. (2021); Shu et al. (2022); Gu et al. (2023) as well as LLM few-shot approaches Li et al. (2023); Nie et al. (2024); Shu and Yu (2024). These require significant volumes of labeled data to train or select the most appropriate few shots.
Also in the in-domain setting, unanswerability has been studied for KBQA Patidar et al. (2023); Faldu et al. (2024).  Patidar et al. (2023) created the GrailQAbility dataset with different categories of unanswerability, and showed that answerable-only models, when adapted superficially for unaswerability, are inadequate in different ways. RetinaQA Faldu et al. (2024) is the SoTA model for unaswerability. It learns to discriminate between valid logical forms and those rendered invalid due to schema-level incompleteness, and constructs logical forms top-down using sketch-filling to handle data-level incompleteness. However, this also requires large volumes of training data.
In the transfer setting for KBQA, one formulation makes use of large volumes of unlabeled training data in the target Cao et al. (2022); Ravishankar et al. (2022). More recently, FuSIC-KBQA has addressed the few-shot transfer setting with only a small volume of labeled training examples in the target Patidar et al. (2024). This uses a retrieve-then-generate framework using one or more supervised retrievers and LLM few-shot in-context learning based generators. The LLM generation also goes through multiple iterations of execution-guided feedback until a non-empty answer is obtained. However, the transfer setting has only considered answerable questions.
LLM prompting techniques have been used to address unanswerability outside of KBQA Slobodkin et al. (2023), but without feedback for iterative generation. Other approaches  Shinn et al. (2023); Chen et al. (2023b) use execution based refinement for programs but do not involve unanswerability.
In the prompting literature, self-consistency marginalizes over multiple reasoning paths to arrive at a consensus answer Wang et al. (2023); Chen et al. (2023a). This works only when significant probability mass accumulates for at least one answer, which is selected as the winner. This breaks down under unaswerability, when reasoning paths to all non-empty answers have low probability.
3 Problem and Proposed Approach
Figure 1: Illustration of FUn-FuSIC for a specific question when executed over three different KBs. The question is answerable for KB3, but unanswerable for KB1 and KB2, due to schema incompleteness and data incompleteness respectively. The iterations as shown as i=1,2,3,4, The consistency block is below these. Colored boxes show different checks. Additions to the set of candidate logical forms are shown in the gray columns.
We first briefly cover background on KBQA, unaswerability and few-shot KBQA. More details are in the Appendix (Sec. A.1). We then define the few-shot setting with unanswerability and our proposed approach.
3.1 Few-shot Transfer Learning for KBQA with Unanswerability
A Knowledge Base $G$ consists of a schema with data stored under it. The schema consists of entity types $T$ and binary relations $R$ defined over pairs of types. The data consists of entities $E$ as instances of types $T$, and triples or facts $F\subseteq E\times R\times E$. We are given a target Knowledge Base $G^{t}$ and a natural language question $q^{t}$, and the goal is to generate a structured query or a logical form $l^{t}$, which when executed over $G^{t}$ returns an answer set $a^{t}$. We are provided with target few-shots $D^{t}$ containing tens of labeled training examples of questions and logical forms. A related source domain has a source knowledge base $G^{s}$ (with its own types, relations, entities and facts), and a larger source training set $D^{s}$ with thousands of labeled training examples. The source and target KBs and the distributions of questions and logical forms can be very different.
In general, a KB may have both answerable and unanswerable questions Patidar et al. (2023). A question $q$ is answerable for a KB $G$ if it has a corresponding logical form which when executed over $G$ returns a non-empty answer $A$. In contrast, a question is unanswerable for $G$, if it either (a) does not have a valid logical form for $G$ (schema-level unanswerability), or (b) it has a valid logical form for $G$, but which on executing on $G$ returns an empty answer (data-level unanswerability). Schema-level unaswerability arises due to missing types (classes) and relations. On the other hand, data-level unaswerability arises due to missing entities and facts. Absence of any mentioned question entity, however, invalidates the logical form. More details are in the Appendix (Sec.  A.1.3).
In KBQA with unanswerability, given a question $q$, the model needs to output (a) a logical form $l$ and a non-empty answer A for answerable $q$, (b) $l=$ NK (No Knowledge) for schema-level unanswerable $q$, or (c) a valid logical form $l$ and $a=$ NA (No Answer) for data-level unanswerable $q$.
We now propose the problem of few-shot transfer learning for KBQA with unanswerability. Here, the target KB $G^{t}$ is incomplete, and target test questions $q^{t}$ may be answerable or unanswerable due to missing schema or data. The few shot examples $D^{t}$ contain both answerable and unanswerable questions of different types. Since almost all large KBQA datasets contain only answerable questions, we assume that the source training data $D^{s}$ contains only answerable questions. So, in addition to the differences between source and target KBQA tasks in the existing transfer learning setting, the unanswerability gap also needs to be bridged.
3.2 Proposed Approach: FUn-FuSIC
Our proposed architecture FUn-FuSIC builds upon and adapts FuSIC-KBQA for unaswerability. We preserve its retrieve-then-generate framework where the first stage is KB retrieval with one or more supervised retrievers, which are source-trained and then target fine-tuned. The retrieval is LLM reranked to increase target relevance. This forms the input to the generative stage, which prompts an LLM to generate the logical form using the retrieval, either in few-shot or zero-shot fashion. More details of these stages are in the Appendix (Sec. F). On top of this FuSIC-KBQA design, FUn-FuSIC makes two modifications. First, the logical form generation instruction includes the possibility of unanswerability, and the few shots include examples of the two types of unanswerable questions. The detailed prompt is in the Appendix (Sec. G.1). Secondly, the generated logical form is refined using feedback as before, but this is now adapted for unaswerability. In the rest of this section, we describe feedback for unaswerability (FUn) in detail.
We observe that the first prong of our solution – prompting for unanswerability (PUn) – largely biases the LLM towards one type of error. For $l=$ NK questions, LLM with PUn mostly generates incorrect logical forms instead of generating NK. The essence of our second prong, FUn, is to assess the LLM’s confidence about any logical form being the correct one for the question. If this confidence is low, then FUn outputs $l=$ NK and $a=$ NA. If there is sufficient confidence in a specific logical form $\hat{l}$, then FUn outputs $l=\hat{l}$. If $\hat{l}$ upon execution returns a non-empty answer $\hat{a}$, then FUn outputs $a=\hat{a}$, else the output is $a=$ NA.
We illustrate flow of FUn using examples in Fig. 1. The complete algorithm is in the Appendix (Sec. 2). FUn assesses confidence in a specific logical form using a suite of checks – syntactic, semantic, and execution-based. Some of these are performed using rules defined for the KB, while others involving uncertainty are performed using LLM prompting. FUn goes over $n$ iterations. In the $i^{th}$ iteration, it generates a new logical form $l^{(i)}$ based on feedback received from checks in earlier iterations. $l^{(i)}$ goes through a sequence of checks. If it passes all checks, FUn outputs $l=l^{(i)}$. If $l^{(i)}$ fails a certain check, it is rejected. On the other hand, if it fails some combination of uncertain checks, it is added to the candidate logical form set $L$. The error message from each failed check is appended to the prompt for the next iteration.
If no clear winner emerges at the end of $n$ iterations, FUn assesses the LLM’s confidence about $L$ as a whole, and also about each logical form in $L$. For this, we propose self consistency with unanswerability (scUn). Let $L_{ne}\subset L$ indicate candidate logical forms that produce non-empty answers upon KB execution. Let $a^{*}$ denote the most popular answer (ties being broken at random) and $L_{ne}(a^{*})\subset L_{ne}$ denote logical forms that produce $a^{*}$. If $|L_{ne}(a^{*})|>t$, where $t$ is a threshold, scUn decides that $a=a^{*}$. LLM prompting then selects the most suitable candidate $l^{*}$ from $L_{ne}(a^{*})$, and FUn outputs ($l=l^{*}$, $a=a^{*}$).
If no consensus emerges from $L_{ne}$, scUn turns to its complement set, $L_{e}=L-L_{ne}$, having empty answers. These represent candidates that agree on $a=$ NA. If $L_{e}$ is not empty, again LLM prompting identifies the most suitable candidate $l^{*}$ from $L_{e}$, and FUn outputs ($l=l^{*}$, $a=$ NA). If on the other hand, $L_{ne}$ is empty, FUn outputs ($l=$ NK, $a=$ NA).
Next, we discuss the different categories of checks. The high level goal is to prune candidate logical forms without knowledge of the gold logical form, the gold answer or even whether the question is unanswerable or not. The specific prompts for the checks and the detailed structures of the feedback containing the error messages are described in the Appendix (Sec. G.2).
(C1) Syntax Error:
The KB engine returns a syntax error upon executing the logical form. This is also a part of EGF in FuSIC-KBQA. Note that this is a certain check – meaning that any logical form that fails this check cannot, with certainty, be the right logical form.
(C2) KB Inconsistency:
The logical form is inconsistent with the KB schema. This can happen in different ways. (C2a) Incompatibility in types: The logical form contains a variable and a connecting relation whose types are incompatible. This error is also checked in FuSIC-KBQA EGF. (C2b) Schema hallucinations: The logical form contains schema elements (types, relations, entities) not in the target KB. (C2c) Type-casting errors: Literals are not correctly type-cast for the target KB, e.g. numeric literals as float for Freebase. All of these are certain checks, and are implemented using rules defined using the logical form and the KB.
(C3) Question Logical form Disagreement:
The generated logical form is syntactically correct and consistent with the target KB but is not semantically equivalent to the original natural language question. This is not a certain check for two reasons. First, different types of ambiguities may exist in the question, ruling out the existence of a semantically equivalent logical form in a specific KB. For example, the question may ask for a person from a country, when the KB only contains the relations born in and works in between these types. Secondly, this check requires a learning based classifier, inevitably having some probability of error.
FUn performs this equivalence check using a novel multi-stage LLM pipeline. (i) The variable names in the logical form are first naturalized, preserving semantics and considering the natural language question, e.g. by replacing ’?x’ with ’?actor’. (ii) The naturalized logical form is back-translated into a natural language question. (iii) The back-translated natural language questions is checked for semantic equivalence with the original natural language question. The first two steps are performed using zero-shot prompting, while the last is performed using few-shots constructed using the target few-shots $D^{t}$. More details about few-shot construction are in the Appendix (Sec. G.3).
(C4) Answer Incompatibility:
If the logical form is syntactically correct and consistent with the target KB, it returns some answer upon execution. This answer may not be compatible with the original question. (C4a) The entity returned as the answer is already contained in the logical form (and therefore mentioned in the question). (C4b) The answer contains at least one entity with an intermediate (compound value) type, and therefore does not correspond to any real-world entity. (C4c) The returned answer is empty. All of these checks can be implemented using rules defined over the logical form, the answer and the KB. Note that while the first two are certain checks, the last is not. In the presence of unaswerability, a question may have $a=$ NA as the gold standard.
We now discuss how FUn identifies candidate logical forms based on these checks and assesses the LLM’s confidence about these.
Identifying Candidate Logical Forms:
In the absence of a clear winner (i.e., a logical form that passes all checks), FUn constructs a candidate set $L$ containing logical forms that are flawed, but not fatally so. Logical forms that fail C1, C2, C4a, or C4b are discarded. Those that fail C3 but pass C4c (have non empty answer but disagree with the question), or fail C4c but pass C3 (have empty answer but are equivalent to the original question) are included in $L$.
Self Consistency for Unanswerability:
Consistency breaks down when all the reasoning paths have low probability, leading to a diverse set of answers with little consensus. For our candidate set $L$ with potentially flawed logical forms, the most popular answer $a^{*}$ may have very few votes. Self consistency for unanswerability (scUn) returns the most popular answer $a^{*}$, only if it receives at least $t$ votes out of $n$. We use $t=\frac{n}{2}$. Otherwise, it returns no consistency.
When there is consistency, scUn outputs the most consistent answer $a^{*}$, empty or non-empty. However, we additionally need to identify the most likely logical form with answer $a^{*}$. When $L(a^{*})\subset L$ has more than one logical form, we use LLM prompting to select the most likely $l^{*}\in L(a^{*})$. Details are in the Appendix.(Sec. A.2)
Adapting FUn for Answerable KBQA:
Though motivated by unanswerability, the core idea of iterative generation of candidate logical forms with feedback from error checks, followed by consistency check is applicable more broadly. To adapt this for answerable KBQA, we make the following simple modifications. (a) Logical forms that fail C4c (empty answer) are not included in the candidate set $L$. (b) Instead of scUn, we fall back on regular self consistency for $L$, so that a consistent answer is always returned. These imply that ($l=l^{*}$, $a=a^{*}$) is the only possible outcome.
4 Experiments
Model WebQSP $\rightarrow$ GrailQAbility WebQSP $\rightarrow$ GraphQAbility
Overall Answerable Unanswerable Overall Answerable Unanswerable
F1(L) EM-s F1 EM-s F1(L) F1(R) EM-s F1(L) EM-s F1 EM-s F1(L) F1(R) EM-s
RetinaQA 58.4 42.2 28.7 26.0 88.0 84.8 58.4 49.7 35.8 18.7 15.2 80.7 78.7 56.4
FuSIC-U 76.6 48.2 67.5 59.2 85.6 80.4 37.2 67.5 34.8 49.3 40.0 85.7 82.8 29.6
FUn-FuSIC 76.1 60.2 67.1 61.2 85.1 80.0 59.2 70.0 53.8 50.7 42.8 89.2 86.5 64.8
Table 1: Performance of different models on two datasets for few-shot KBQA transfer with unanswerability.
We now present the experimental evaluation of FUn-FuSIC. First, we aim to address the following research questions for the few-shot transfer setting for KBQA with unanswerability. (R1) How does FUn-FuSIC compare against SoTA models for KBQA with unanswerability, adapted for the few-shot setting? (R2) How does FUn-FuSIC perform across different categories of unanswerability? (R3) To what extent do the different components of FUn-FuSIC contribute to its performance? In addition, we ask the following research question for the answerable KBQA setting. (R4) How does FUn-FuSIC compare against SoTA few-shot transfer (answerable) KBQA models?
4.1 Experimental Setup
Datasets:
[…]
We create our target test sets by selecting 250 answerable and 250 unanswerable questions uniformly at random from the GrailQAbility and GraphQAbility test sets. Additionally, we create the target few shots by selecting 50 answerable and 50 unanswerable questions uniformly at random from the GrailQAbility dev set and the GraphQAbility train set respectively. Of the answerable questions, schema level unanswerable forms 67% in GrailQAbility and 75% in GraphQAbility. The source WebQSP training set has 2858 questions, all of which are answerable.
Model WebQSP $\rightarrow$ GrailQAbility WebQSP $\rightarrow$ GraphQAbility
Schema Level Data Level Schema Level Data Level
F1(L) F1(R) EM-s F1(L) F1(R) EM-s F1(L) F1(R) EM-s F1(L) F1(R) EM-s
RetinaQA 94.1 90.9 79.4 76.3 72.9 14.1 83.2 82.0 72.3 73.7 72.7 12.1
FuSIC-U 85.4 80.6 30.9 86.0 80.0 49.4 86.6 82.6 19.0 83.3 83.3 51.5
FUn-FuSIC 85.8 81.2 70.9 83.8 77.6 36.5 92.4 87.5 75.6 80.3 80.3 34.8
Table 2: Performance of various models for different categories of unanswerable questions in two datasets.
Models for comparison:
There are no existing few-shot transfer KBQA models for unanswerability. RetinaQA Faldu et al. (2024) is the SoTA in-domain model for KBQA with unanswerability. It has been shown to significantly outperform various unanswerability adaptations of all other SoTA in-domain KBQA models, such as RnG-KBQA Ye et al. (2022), TIARA Shu et al. (2022), and Pangu Gu et al. (2023), across different categories of unanswerability. We use the available code for RetinaQA. [^2] https://github.com/dair-iitd/RetinaQA We train all components on WebQSP, using as dev set the corresponding target domain dev set. In the absence of unanswerable questions for training, RetinaQA uses a threshold fine-tuned on a dev set to detect schema-level unanswerability. We again use the target dev sets for this. More details are in the Appendix (Sec. I).
FuSIC-KBQA Patidar et al. (2024) is the SoTA model for few-transfer for KBQA with only answerable questions. It comprehensively outperforms a host of supervised and LLM-equipped KBQA models, such as TIARA Shu et al. (2022), Pangu Gu et al. (2023), KB-Binder Li et al. (2023), gf-LLM Shu and Yu (2024), adapted for few-shot transfer. The logical form generation prompt of FuSIC-KBQA is adapted for unanswerability in the instruction and also the few-shots, which now contain unanswerable exemplars. Specifically, we use 5 shots, of which 2 are answerable, 1 is data-related unanswerable, 1 is schema-related unanswerable, and 1 has missing mentioned entity. This adaption also eliminates execution-guided feedback (EGF) from FuSIC-KBQA, which, being based entirely on empty answer check, is unusable for unanswerability. We name this model FuSIC-KBQA-U. Observe that FuSIC-KBQA-U can also be seen as an ablation of FUn-FuSIC, where FUn is removed.
Recall that FUn-FuSIC has two mechanisms for addressing unanswerability via LLM prompting. The first is providing few-shots with unanswerability for logical form generation (PUn). The second is using $n$ iterations of FUn. For a fair comparison, all LLM equipped models should have the same LLM “budget” – equivalently, cumulative prompt size. For this reason, we use FUn-FuSIC with zero-shot generation and 4 iterations of FUn, which roughly equals the prompt length of 5-shot FUn-FuSIC-U.
We use gpt-4-0613 as the LLM for both FUn-FuSIC and FuSIC-KBQA-U. The FUn-FuSIC framework allows flexible use of supervised retrievers. For meaningful comparison, we adapt and use RetinaQA as the retriever for both FUn-FuSIC and FuSIC-KBQA-U. More details about the retriever along with details of compute infrastructure are in the Appendix (Sec. H).
Evaluation Measures:
Following existing literature Patidar et al. (2023); Faldu et al. (2024), we evaluate answers using regular and lenient F1 (F1(R) and F1(L)). F1(L) additionally gives credit for the original answer for the complete KB.
For logical form evaluation, the existing evaluation measure (EM) Ye et al. (2022) works only for s-expression as the language. Since FuSIC-KBQA-U and FUn-FuSIC output logical forms in SPARQL and RetinaQA in s-expression, we propose a new measure (EM-s) that works for both.
EM-s considers two logical forms to be equivalent if these contain identical sets of relations and entities, and additionally return identical answer sets upon KB execution. Since this is a necessary but not sufficient condition for logical form equivalence, we compared EM-s with EM, where both are applicable, and found $>98\%$ agreement.
More details are in the Appendix (Sec. C).
Answerable Unanswerable
F1 EM-s F1(L) F1(R) EM-s
FUn-FuSIC 57.0 46.0 92.0 90.0 62.0
sc 63.0 52.0 66.0 64.0 36.0
w/o syntax 57.0 46.0 92.0 90.0 62.0
w/o kb-inc 51.7 42.0 92.0 88.0 44.0
w/o q-lf 47.7 38.0 55.8 48.0 10.0
w/o ans-inc 55.0 44.0 92.0 90.0 62.0
Table 3: Ablation performance of FUn-FuSIC (removing individual components with replacement) on WebQSP $\rightarrow$ GraphQAbility. sc denotes replacing self consistency for unanswerability with self consistency. Other rows remove checks for syntax error (syntax), KB inconsistency (kb-inc), question logical form disagreement (q-lf) and answer incompatibility (ans-inc).
4.2 Unanswerability Setting
We first address research question R1. Performances of the different models for the few-shot transfer setting with unanswerability are recorded in Tab. 1. We make the following observations. (a) For answerable questions, FuSIC-U and FUn-FuSIC significantly outperform RetinaQA. Across the two datasets, FUn-FuSIC mostly performs better than FuSIC-U (both F1 and EM-s). (b) For unanswerable questions, FuSIC-U performs well in terms of the F1 scores, but not EM-s, indicating empty answers with incorrect logical forms. RetinaQA performs well in terms of F1 scores and EM-s. FUn-FuSIC performs slightly worse than RetinaQA for WebQSP $\rightarrow$ GrailQAbility, but significantly better for WebQSP $\rightarrow$ GraphQAbility. (c) Overall, FUn-FuSIC significantly outperforms both baselines in terms of EM-s, and is slightly below FuSIC-U in terms of F1(L) for WebQSP $\rightarrow$ GrailQAbility. This establishes usefulness of FUn for few-shot transfer KBQA with unanswerability.
We next briefly address research question R2.
Performance of different models for different categories of unanswerability are recorded in Tab. 2. While all models struggle with data-level unanswerability (EM-s), RetinaQA performs the worst and FuSIC-U is the best. For schema-level unanswerability, FUn-FuSIC is the best in one dataset, while RetinaQA is the best in the other. This indicates that more work is needed for robustness across unanswerability categories, and simultaneous use of PUn (prompting) and FUn (feedback) is the most promising approach.
We next address research question R3. Tab. 3 records performance for different ablations of FUn-FuSIC. The biggest benefit, for both answerable and unanswerable questions, comes from question logical form disagreement feedback. Feedback on KB inconsistency and answer incompatibility also make significant contributions to the performance.
Model WebQSP $\rightarrow$ WebQSP $\rightarrow$
GrailQA-Tech GraphQA-Pop
FuSIC 70.8 52.3
FUn-FuSIC(sc) 73.6 67.0
FuSIC-U 62.6 43.4
FUn-FuSIC(scUn) 71.2 65.0
Table 4: Performance using F1 of different models for few-shot KBQA transfer with only answerable questions. The models in the first two rows assume knowledge of answerability of all questions, while the those in the next two do not.
4.3 Answerable Setting
We now address answerable-only KBQA and research question R4. We use two of the datasets used in existing literature for this setting, including the hardest one (WebQSP $\rightarrow$ GraphQA-Pop) Patidar et al. (2024) [^3] https://github.com/dair-iitd/FuSIC-KBQA/. For consistency with existing literature, here we use TIARA Shu et al. (2022) as the retriever for all models.
This setting can be evaluated in two ways. In the first setting, the models are aware that all questions are answerable. This is the setting studied in the KBQA literature, and FuSIC-KBQA is the established SoTA model outperforming a host of supervised and LLM-based models adapted for the task Patidar et al. (2024). We appropriately adapt FUn-FuSIC as discussed earlier in Sec. 3.2.
The first two rows in Tab. 4 record performance for this setting. We see that FUn-FuSIC with self-consistency significantly outperforms FuSIC-KBQA on both datasets, creating a new SoTA for this setting. This establishes the usefulness of the broader idea of iterative generation of logical forms, with feedback from a diverse suite of error checks.
In the more practical second setting, which has not been studied before, although all questions are answerable, the models are not aware of this and make predictions assuming unanswerability. For this we evaluate FuSIC-U and FUn-FuSIC as in Sec. 4.2. The bottom two rows of Tab. 4 record their performance. We see that FUn-FuSIC outperforms FuSIC-KBQA by a very large margin. This further establishes the usefulness of scUn when guarantees about answerability do not exist.
4.4 Error Analysis
We have performed detailed error analysis for FUn-FuSIC. For WebQSP $\rightarrow$ GraphQAbility, we analyzed questions with EM-s $<1$ (46.2% of questions). There are two high-level error categories. (E1) Retrieval Error: At least one necessary KB element (class, relation, entity) is missing in the retrieved KB context (23.4% of questions). (E2) Generation Error: Despite correct retrieval, the generated logical form is incorrect (22.8% of questions). Generation errors can be further sub-categorized. (E2a) The gold logical form is NK, but some (non-NK) logical form is predicted (8.4% of questions). (E2b) A (non-NK) gold logical form exists, but NK is predicted, (4.6% of questions). (E2c) The gold and the predicted logical forms are non-NK, but different (9.8% of questions). We found three main causes for generation errors. (1) Some questions are inherently ambiguous, admitting multiple valid logical forms in a complete KB. Incompleteness eliminates one of these, so that the gold logical form is NK, while the model identifies the other still existing non-NK logical form. (2) FUn is unable to generate the gold logical form within its maximum number of iterations ($n=4$). (3) FUn generates the gold logical form but it is rejected incorrectly by the question logical form disagreement checker.
5 Conclusions
For robust, low-resource KBQA, we have proposed the practical task of few-shot transfer learning with unanswerability. We have proposed FUn-FuSIC that adapts the existing SoTA model that assumes answerability by (i) candidate logical form generation by iterative LLM prompting with feedback from diverse error checks – including a novel back-translation based equivalence check, and (ii) assessing confidence in the candidates by adapting self consistency to detect answerability and different categories of unanswerability. Using newly created datasets, we show that FUn-FuSIC outperforms adaptations of SoTA models for this setting, and also for answerable KBQA. Our error analysis suggests that performing well across all categories of unanswerability in the few-shot transfer setting is still a challenge and should be a focus of further research.
6 Limitations
Since LLM inference involves randomness, ideally experiments should be repeated for multiple runs and results should report averages and error bars. Unfortunately, we were not able to do this due to the prohibitive cost of GPT-4, and our results are based on single runs.
While GPT-4 is currently the best performing LLM, it is proprietary as well as expensive. Ideally, evaluation should involve open-source freely accessible LLMs as well. We expect performance of all LLM-based approaches to drop when GPT-4 is replaced by an open LLM. Earlier research has shown that models using Mistral are still able to outperform fully supervised models for answerable few-shot transfer Patidar et al. (2024). Whether this trend holds for the unanswerable setting is an open question. That said, following current trends, we expect the ability of open LLMs to steadily improve in the coming years.
7 Risks
At the highest level, our work reduces risk compared to existing KBQA systems, which when improperly adapted in a low-resource setting, incorrectly answer unanswerable questions, without admitting lack of knowledge. However, can incorrectly inferring unaswerability, citing lack of knowledge when the knowledge is in fact available, generate a new category of risk? While we cannot imagine such a risk at the present time, this may require more thought. In any case, KBQA models for unanswerability should strive to minimize this type of error, along with the other types.
Appendix A Appendix
A.1 KBQA Problem Formulations
We begin by defining the few-shot transfer learning setting for Knowledge Base Question Answering (KBQA) with unanswerability, and then present our solution for it.
A.1.1 KBQA:
A Knowledge Base $G$ consists of a schema with data stored under it. The schema consists of entity types $T$ and binary relations $R$ defined over pairs of types. The data consists of entities $E$ as instances of types $T$, and triples or facts $F\subseteq E\times R\times E$. We are given a target Knowledge Base $G^{t}$ (consisting of entity types $T^{t}$, relations $R^{t}$, entities $E^{t}$ and facts $F^{t}$) and a natural language question $q^{t}$, and the goal is to generate a structured query or a logical form $l^{t}$, which when executed over $G^{t}$ returns a non-empty answer $A^{t}$ for the question $q^{t}$.
A.1.2 Few Shot Transfer Learning for KBQA:
In few-shot transfer learning, we are provided with target few-shots $D^{t}$ containing tens of labeled training examples of questions and logical forms in the target domain. In addition, we are given access to a related source domain. This has a source knowledge base $G^{s}$ consisting of its own types $T^{s}$, relations $R^{s}$, entities $E^{s}$ and facts $F^{s}$, and a much larger source training set $D^{s}$ containing thousands of labeled training examples of questions and corresponding logical forms. The source and target tasks may differ significantly. First, the data and schema of the knowledge bases $G^{t}$ and $G^{s}$ and the domains they cover may be different. Secondly, the distributions of questions and logical forms defined over the KBs may be different in $D^{t}$ and $D^{s}$.
A.1.3 KBQA with unanswerability:
A natural language question q is said to be answerable for a Knowledge Base $G^{t}$ if it has a corresponding logical form $l$ which when executed over $G^{t}$ returns a non-empty answer $A$. In contrast, a question $q$ is unanswerable for $G^{t}$, if it either (a) does not have a corresponding logical form that is valid for $G^{t}$, or (b) it has a valid logical form $l$ for $G^{t}$, but which on executing returns an empty answer. The first case indicates that $G^{t}$ is missing some schema element necessary for capturing the semantics for $q$. The second case indicates that the types $T^{s}$, relations $R^{s}$ is sufficient for $q$, but $G^{t}$ is missing some necessary data elements for answering it. In the KBQA with unanswerability task, given a question $q$, if it is answerable, the model needs to output the corresponding logical form $l$. and the non-empty answer $A$ entailed by it, and if it is unanswerable, the model either needs to output $NK$ (meaning No Knowledge) for the logical form, or a valid logical form $l$ with $NA$ (meaning No Answer) as the answer.
Different Types of Unanswerability:
Unanswerable questions in KBQA can be categorized into (a) Schema Level Unanswerability : the question does not have a corresponding logical form that is valid for the KB , (b) Data level unanswerability: it has a valid logical form $l$ for the KB, but which on executing returns an empty answer. Schema level unanswerable questions can further be categorized into (1) Missing Class: The class/type required to construct the logical form is not defined for the KB, (2) Missing Relation: The relation required to construct the logical form is not defined for the KB, (3) Missing Entity: The topic entity specified in the question is missing from the KB. Data level unanswerable questions can be categorized into (1) Missing entity: all classes and relations required to construct the logical form are present in the KB , but there exists no path from the topic entity node to the answer node in the KB due to missing intermediated entities (2) Missing Fact: all classes, relations and entities required to answer the question are present in the KB. However, the (subject, relation, object) path is not connected in the KB.
A.1.4 Few-shot Transfer Learning for KBQA with Unanswerability:
This setting is a generalisation of few shot transfer learning for KBQA, wherein we no longer assume that the test question is answerable with respect to the Knowledge Base $G^{t}$. The training set $D^{s}$ contains thousands of answerable questions along with their corresponding logical form. The target few shots $D^{t}$ contain both answerable as well as unanswerable questions, along with their corresponding logical forms, if they exist.
A.2 Algorithm
Figure 2: Feedback Algorithm for Robustness in KBQA
1:Given: KB-Engine, KB-Ontology, qn $\triangleright$ qn is natural language query
[…]
3:$L\leftarrow[]$
4:for $i=0$ to $N$ do
5:     $lf\leftarrow LLM(prompts)$
6:     $syntax\_error\leftarrow KB-Engine(lf)$
7:     if $syntax\_error\neq\text{null}$ then
8:         $syntax\_prompt\leftarrow SynPrompt(lf,syntax\_error)$
9:         $prompts.append(syntax\_prompt)$
10:         continue
11:     end if
12:     $kb\_consistency\_error\leftarrow CheckTypes(lf,KB-Ontology)$
13:     if $kb\_consistency\_error\neq\text{null}$ then
14:         $kb\_consistency\_prompt\leftarrow LFPrompt(lf,kb\_consistency\_error)$
15:         $prompts.append(kb\_consistency\_prompt)$
16:         continue
17:     end if
18:     $qlf\_error\leftarrow CheckMeaning(lf,qn)$
19:     $ans\_consistency\_error\leftarrow CheckAnswer(KB-Engine(lf),lf.nodes)$
20:     if $qlf\_error\neq\text{null}$ then
21:         $qlf\_prompt\leftarrow NLPrompt(lf,qlf\_error)$
22:         $prompts.append(qlf\_prompt)$
23:         if $ans\_consistency\_error$ = null then
24:              $L.append(lf)$
25:         end if
26:         continue
27:     end if
28:     if $ans\_consistency\_error\neq\text{null}$ then
29:         $ans\_consistency\_prompt\leftarrow EGF(lf,egf\_error)$
30:         $prompts.append(ans\_consistency\_prompt)$
[…]
32:              $L.append(lf)$
33:         end if
34:         continue
35:     end if
36:     return $lf$triangleright all checks passed
37:end for
38$\triangleright$ despite N-tries, no lf passes
39:$L_{ne}\leftarrow[l\text{ for }l\text{ in }L\text{ if }KB-Engine(l)\neq\text{null}]$
40:$L_{e}\leftarrow[l\text{ for }l\text{ in }L\text{ if }KB-Engine(l)=\text{null}]$
41:$poss\_answers\leftarrow[KB-Engine(lf)\text{ for }lf\text{ in }L_{ne}]$
42:if $scUn(poss\_answers)\neq\text{null}$ then $\triangleright$ No Answer-aware-self-consistency
43:     return $SelectBestLF(qn,[lf\text{ for }lf\text{ in }L_{ne}\text{ if }KB-Engine(lf)==scUn(poss\_answers)])$
44:else if $is\_unanswerable$ then
45:     if $L_{e}\neq[]$ then
46:         return $SelectBestLF(qn,L_{e})$
47:     end if
48:     return NK
49:else
50:     return $SelectBestLF(qn,[lf\text{ for }lf\text{ in }L_{ne}\text{ if }KB-Engine(lf)==sc(poss\_answers)])$triangleright self-consistency, with tie breaking
51:end if
Algorithm 1 Self-Consistency Algorithm
1:Given: $possible\_answers$
2:Filter out null values from possible_answers:
3:$possible\_answers\leftarrow[p\text{ for }p\text{ in }possible\_answers\text{ if }p\neq\text{null}]$
4:if $possible\_answers=[]$ then
5:     return null
6:end if
7:$MajVote\leftarrow MaxFreqItem(possible\_answers)$
8:return $MajVote$
Algorithm 2 Self-Consistency with Unanswerability Algorithm
1:Given: $possible\_answers$
2:$MajVote\leftarrow self-consistency(possible\_answers)$
3:if $MajVote=\text{null}$ then
4:     return null
5:end if
6:$count\leftarrow Cnt(MajVote,possible\_answers)$
7:if $count>\frac{\text{len}(possible\_answers)}{2}$ then
8:     return $MajVote$
9:else
10:     return null
11:end if
A.3 Working Example of The Algorithm
Question: Where all were the Olympics held before 2024?
Logical Form 1 ?y{ ?x olympics.venue ?y
    ?y olympics.year ?z
    FILTER(?z = 2024)}
Feedback 1 Type: Logical Form Semantic Error Description: ?y is assigned types [’venue’, ’olympics’] by relations [’olympics.venue’, ’olympics.year’]. These types are mutually incompatible.
Logical Form 2 ?y{ ?x olympics.venue ?y
    ?x olympics.year ?z
    FILTER(?z = 2024)}
Feedback 2 Type: Natural Language Semantic Error Description: The question originally asked returns all places where the Olympics were held before 2024. However, the question we answer returns the places where the Olympics were held in 2024.
Logical Form 3 ?y{ ?x olympics.venue ?y
    ?x olympics.year ?z
    FILTER(?z < 2024)}
Feedback 3 Status: Passes All Checks Details: lf3 satisfies syntax, logical form semantic, natural language semantic and execution checks.
Appendix B Ablations
We perform ablations of our best performing model - FUn-FuSIC by removing different types of feedback from the model. We report performance on 100 questions(50 answerable + 50 unanswerable) sampled from the WebQSP $\rightarrow$ GraphQAbility dataset. Refer to 5
Ablation WebQSP $\rightarrow$ GraphQAbility
Overall Ans Unans
F1(L) SM F1 SM F1(L) F1(R) SM
FUn-FuSIC 74.5 54.0 57.0 46.0 92.0 90.0 62.0
sc 64.5 44 63.0 52.0 66.0 64.0 36.0
w/o syntax 74.0 54.0 57.0 46.0 92.0 90.0 62.0
w/o kb-inc 71.8 43.0 51.7 42.0 92.0 88.0 44.0
w/o qf-lf 51.7 24.0 47.7 38.0 55.8 48.0 10.0
w/o ans-inc 73.5 53.0 55.0 44.0 92.0 90.0 62.0
Table 5: Detailed ablation Performance of FUn-FuSIC on WebQSP $\rightarrow$ GraphQAbility. sc denotes replacing self consistency for unaswerability with standard self consistency, while other rows indicate removing check for syntax error (syntax), KB inconsistency (kb-inc), question logical form disagreement (q-lf) and answer incompatibility (ans-inc).
Appendix C EM-s: EM for sparql
As has been observed in Patidar et al. (2023), answer evaluation by itself is not a robust measure for evaluation of KBQA models when the dataset contains unanswerability. Traditional KBQA models that generate s-expressions can be evaluated using EM, which checks for logical form equivalence between two logical forms, since it is possible to compare equivalence between two s-expressions efficiently. However, FUn-FuSIC generates sparql queries instead. Directly comparing program equivalence between two sparql queries is an undecideable problem  [^4] https://users.dcc.uchile.cl/ cgutierr/papers/expPowSPARQL.pdf. Patidar et al. (2024) suggests a semi-automatic strategy for comparison of sparql queries. We propose a completely automatic metric for sparql equivalence check. Two sparql queries are equivalent by the EM-s check if (a) the relations occuring in the two queries are same. (b) the entities occuring in the two queries are the same (c) the answer set obtained by executing the queries over the KB are the same. Note that the EM-s check is necessary, but not sufficient for two sparql queries to be equivalent.
Since these are a necessary but not sufficient condition for logical form equivalence, we compared EM-s with EM, where both are applicable and found $>98\%$ agreement.
Appendix D Performance across different categories of unanswerability
We dive deeper into the different categories of unanswerability, as explained in Patidar et al. (2023). There exist 2 broad categories of unanswerability- schema level unanswerability(absence of knowledge in terms of KB ontology or entities required to construct the logical form) and data level unanswerability(absence of facts or intermediate entities of the logical form path on the KB).
We expect that (a) due to the poor ability of supervised models to generalize in transfer learning settings, RetinaQA will be struggle to generate correct logical forms for data level unanswerable questions (b) due to the strong generalization ability of FuSIC-KBQA, it should be able to perform well for data level unanswerable questions. However, since it is biased to return incorrect logical forms instead of abstaining from returning a logical form, it will perform poorly at identifying schema level unanswerable questions. (c) FUn-FuSIC should be able to maintain the performance of FuSIC-KBQA on data level unanswerable questions to a large extent, while significantly improving the performance on schema level unanswerable questions.
Based on performance on the WebQSP $\rightarrow$ GrailQAbility and WebQSP $\rightarrow$ GraphQAbility datasets, we validate the that the trends are indeed as expected.
Appendix E Cost Analysis
FuSIC-KBQA, as well as the adapted versions of FuSIC-KBQA, such as U-FuSIC and FUn-FuSIC rerank the classes, relations and paths. The total cost for reranking for one question is \$0.16.
The cost for generation of logical form from a prompt with 5 in-context examples is \$0.16.
Thus, the approximate cost for inference of one question by U-FuSIC is \$0.32.
The cost for generation of logical form from a prompt with 0 in-context examples is \$0.04. The cost of checking whether two natural language questions are equivalent or not, using few-shot exemplars and chain of thought prompting is also \$0.04.
The approximate cost of inference of one question by FUn-FuSIC varies between \$0.24 and \$0.48. The average cost over 50 randomly sampled questions from the test set is around \$0.34.
Hence, the two models are comparable in terms of cost.
Appendix F FuSIC-KBQA: High Level Overview
Our proposed approach FUn-FuSIC builds upon the the base architecture of FuSIC-KBQA. FuSIC-KBQA has a three step pipeline: (a) Supervised Retrieval: a supervised retriever, trained on the source domain and optionally fine-tuned on the target domain is used to obtain the top-100 classes, relations and paths that are relevant to the question asked, (b) LLM ReRanking: an LLM is used to select the top-10 classes, top-10 relations and top-5 paths among the top-100 elements as ranked by the supervised retriever, (c) LLM Generation: We provide the top-10 classes, top-10 relations and top-5 paths along with few-shot exemplars to generate the sparql query.
Appendix G Prompts
G.1 PUn prompt
Header Prompt Translate the following question to sparql for Freebase based on the candidate sparql, candidate entities, candidate relations and candidate entity types which are separated by "|" respectively. Please do not include any other relations, entities and entity types. Your final sparql can have three scenarios: 1. When you need to just pick from candidate sparql. 2. When you need to extend one of candidate sparql using the candidate relations and entity types. 3. When you will generate a new sparql only using the candidate entities, relations and entity types.For entity type check please use this relation "type.object.type".Do not use entity names in the query. Use specified mids. If it is impossible to construct a query using the provided candidate relations or types, return "NK"Make sure that the original question can be regenerated only using the identified entity types, specific entities and relations.
NK exemplar Question: the tv episode segments spam fall under what subject? Candidate entities: spam m.04vbm Candidate paths: SELECT DISTINCT ?xWHERE ?x0 ns:tv.tv_segment_performance.segment ns:m.04vbm .?x0 ns:tv.tv_segment_performance.segment ?x .?x ns:type.object.type ns:tv.tv_episode_segment . | ... Candidate entity types: tv.tv_series_episode| tv.tv_episode_segment | ... Candidate relations: tv.tv_series_episode.segments (type:tv.tv_series_episode R type:tv.tv_episode_segment)| tv.tv_subject.tv_programs (type:tv.tv_subject R type:tv.tv_program)|... sparql:NK
Question Prompt Question: which school newspaper deals with the same subject as the onion? Candidate entities: the onion m.0hpsvmv Candidate paths: SELECT DISTINCT ?xWHERE ns:m.0hpsvmv ns:book.newspaper.circulation_areas ?x0 .?x0 ns:periodicals.newspapers ?x .?x ns:type.object.type ns:book.newspaper . |... Candidate entity types: education.school_newspaper| type:book.newspaper... Candidate relations: education.school_newspaper.school (type:education.school_newspaper R type:education.educational_institution) | book.newspaper_issue.newspaper (type:book.newspaper_issue R type:book.newspaper)|... sparql:
G.2 FUN prompt
syntax error feedback Correct the syntax of the following sparql query. Return ONLY the corrected sparql query without any explanation sparql: SELECT ?x AND ?y ... Virtuoso error: word AND not defined
Target KB Disagreement feedback The generated sparql has a semantic issue warning: The types of relations don’t match for variable ?x in the query. The assigned relation types by [’computer.computer_emulator.computer’, ’type.object.type computer.computer_peripheral’] are [’computer.computer’, ’computer.computer_peripheral’]. These types are mutually incompatible... Please generate again a different executable sparql using the same context and constraints. DO NOT APOLOGIZE - just return the best you can try.
Question Logical form disagreement feedback The question that you answer is NOT same as what you’ve been asked for! You have answered the question "Which opera productions has Gino Marinuzzi conducted?" but you were asked to answer "what is the name of the premiere opera production conducted by gino marinuzzi?". Please generate again a different executable sparql using the relations, classes and entities provided earlier. DO NOT APOLOGIZE - just return the best you can try.
The next three prompts fall under the answer incompatibility feedback
empty answer feedback The generated sparql gives an empty answer when executed on freebase KG, Please generate again a different executable sparql using the same context and constraints.
intermediate node feedback The generated sparql returns an intermediate type node when executed on the freebase KG. Maybe the answer node is an adjacent node to what we currently query for. Please generate again a different executable sparql using the same context and constraints.
answer node in question feedback The logical form upon execution returns International System of Units, which is not answering the question. Please reconstruct the query using same context and constraints.
G.3 Question Logical Form Agreement Prompt
The few shots provided for verifying question logical form agreement come from the few shots provided in the target domain $D^{t}$. We obtain positive samples from the dataset $D^{t}$ directly, using the questions and gold logical forms. For obtaining negative samples, we perform zero-shot FuSIC-KBQA inference over the dataset $D^{t}$. Then we consider those questions for which the predicted logical form is different from the gold logical form.
Firstly, we perform back-translation to obtain natural language question from the logical form
Naturalization of variable names change the sparql query to have variable names representative of what objects they refer to. transform the variable names in this query. Do NOT change the prefix headers and relation names
[…]
We use few-shot LLM prompting to obtain the explanation for why the question and logical form agree or disagree. These few-shots, for obtaining the explanation are dataset independent, and are manually written.
explanation generation prompt Explain why the two questions are different. Question we answer: who all like to eat apple or mango? Question originally asked: what are the people who enjoy both apple and mango? explanation: The question we answer returns people. The question originally asked also returns people. The question we answer finds those people who like eating apple, those people who like eating apple. The question originally asked also finds those people who like eating apple, those people who like eating apple. The question we answer uses logical operator OR. However, the question originally asked uses the logical operator AND Hence, they are different. [total 3 exemplars] Question we answer: Which game engines are successors to the Unreal Engine? Question originally asked: which video game engine’s successor engine is unreal engine? explanation:
Some examples of few shots are provided below-
question logical form agreement check Question we answer: Who are the cricket players who have made exactly 31 stumps in one day internationals? Question originally asked: name the cricket player who has 31 odi stumps. explanation: The question we answer returns cricket players. The question originally asked also returns cricket players. The question we answer finds cricket players who have made exactly 31 stumps in one day internationals. The question originally asked also finds cricket players who have made 31 stumps in one day internationals. Both questions involve no mathematical or logical operators. Hence, they are same. Question we answer: Which game engines are successors to the Unreal Engine? Question originally asked: which video game engine’s successor engine is unreal engine? explanation: The question we answer returns game engines. The question originally asked also returns game engines. The question we answer finds successors to the Unreal Engine. The question originally asked finds the predecessor of the Unreal Engine. The reasoning steps followed by the two questions are different. Hence, they are different. [total 6 exemplars] Question we answer: Which cars drive at a speed of 80? Question originally asked: name the car with driving speed atleast 80? explanation:
G.4 Logical Form Selection Prompt
We store the back-translated natural language questions along with each predicted logical form. Here, we prompt the LLM to select the closest back-translated natural language question to the orginal question.
Best Logical Form Selection Prompt orig_nl_qn = which surf films has sarah finn served as the casting director? 1. pred_nl: Which surfing films has Sarah Finn directed the casting for? 2. pred_nl: Which surfing films has Sarah Finn been the casting director for? of the 2 predicted nl questions, which is closest to the original nl question. Even if none is very close, return the one that is semantically closest? pls explain your answer as well
Appendix H Supervised Models training details
We use Hugging Face  Wolf et al. (2020), PyTorch Paszke et al. (2019) for our experiments and use the Freebase setup specified on github 555 https://github.com/dki-lab/Freebase-Setup . We use NVIDIA A100 GPU with 40 GB GPU memory and 32 GB RAM. For training the discriminator module of RetinaQA, we require 2 GPUs. (1) For the answerable experiments, we use the supervised models as specified in Patidar et al. (2024). (2) For the unanswerability experiments, we train all models from scratch. (a) We use RnG-KBQA entity linker  [^6] https://github.com/salesforce/rng-kbqa/tree/main/GrailQA/entity_linker (BSD 3-Clause License) trained on the answerable subset of GrailQAbility for all our experiments. (b) We train the RnG-KBQA path retriever on answerable subset of WebQSP [^7] https://github.com/salesforce/rng-kbqa/blob/main/WebQSP/scripts/run_ranker.sh (BSD 3-Clause License). The number of training epochs is determined by the performance of the model over the answerable questions in the dev set. (c) We train the TIARA schema retriever on the answerable subset of WebQSP  [^8] https://github.com/microsoft/KC/tree/main/papers/TIARA/src (MIT License) (d) We train the sketch generator and discriminator of RetinaQA on the answerable subset of WebQSP [^9] https://github.com/dair-iitd/RetinaQA .
Appendix I Supervised Models inference details
We use the dev set in RetinaQA, during discriminator inference for (a) determining how to best utilize the candidate paths. We might (i) not provide candidate paths (ii) provide candidate paths in GrailQA format (iii) provide candidate paths in WebQSP format. We select the best alternative based upon the performance of the model over the dev set. For the WebQSP $\rightarrow$ GrailQAbility dataset, we observe (ii) works best, whereas for the WebQSP $\rightarrow$ GraphQAbility dataset, we observe (i) works best. (b) determining the threshold value. RetinaQA applies a threshold on the scores - for a question, if the highest score candidate logical form has a score less than the threshold, the question is labelled as NK. We choose the optimal value of the threshold to maximise the overall EM-s score over the dev set.