### v3

# Iterative Repair with Weak Verifiers for Few-shot Transfer in KBQA with Unanswerability

Abstract
Real-world applications of KBQA require models to detect different types of unanswerable questions with a limited volume of in-domain labeled training data. We propose the novel task of few-shot transfer for KBQA with unanswerable questions. The state-of-the-art KBQA few-shot transfer model (FuSIC-KBQA) uses an iterative repair strategy that assumes that all questions are answerable. As a remedy, we present FUn-FuSIC – a novel solution for our task that extends FuSIC-KBQA with Feedback for Unanswerability (FUn), which is an iterative repair strategy for answerable as well as unanswerable questions. FUn uses feedback from a suite of strong and weak verifiers, and an adaptation of self-consistency for unanswerability for assessing answerability of questions. Our experiments show that FUn-FuSIC significantly outperforms suitable adaptations of multiple LLM-based and supervised SoTA models on our task, while establishing a new SoTA performance for answerable few-shot transfer as well. We have made datasets and other resources publicly available [^1] https://github.com/dair-iitd/FUn-FuSIC
Iterative Repair with Weak Verifiers
for Few-shot Transfer in KBQA with Unanswerability
Riya Sawhney‡,§, Samrat Yadav, Indrajit Bhattacharya, Mausam Indian Institute of Technology, Delhi, §Graviton Research Capital, KnowDis AI riya.sawhney@outlook.com, samratya23@gmail.com, indrajitb@gmail.com, mausam@cse.iitd.ac.in
1 Introduction
The semantic parsing formulation of the Knowledge Base Question Answering (KBQA) task takes as input a Knowledge Base (KB) and a natural language question, and outputs a logical form (or program) that produces the answer upon execution over the KB. KBQA has important real-world applications, which require KBQA systems to be low-resource (i.e., trained only with a few task-specific labeled examples), and robust, specifically able to identify questions that cannot be answered from the KB.
Traditional supervised models (e.g., Ye et al. (2022); Shu et al. (2022); Gu et al. (2023)) and even recent LLM few-shot in-context learning (FS-ICL) architectures Li et al. (2023); Nie et al. (2024) for KBQA fall short in both aspects. Limited recent work has addressed these independently – in-domain methods for KBQA with unanswerability trained with large labeled data Patidar et al. (2023); Faldu et al. (2024), and FuSIC-KBQA for few-shot transfer assuming answerable questions Patidar et al. (2024). No existing single KBQA model simultaneously addresses both desiderata.
In response, we propose the novel task of few-shot transfer learning for KBQA with unanswerability. Specifically, the target domain has only a few labeled examples of answerable and unanswerable questions, while the source domain has thousands of labeled examples, but containing only answerable questions.
For few-shot KBQA transfer, FuSIC-KBQA uses a retrieve-then-generate framework: retrieval of relevant schema and KB snippets followed by an LLM-based generation and a subsequent iterative execution-error-guided repair. Specifically, multiple feedback-guided repair iterations are executed, checking emptiness of answers obtained by executing the generated program as indication of correctness, until a non-empty answer is obtained. This naturally fails when questions are allowed to be unanswerable.
A simple-fix for addressing Unanswerability (FuSIC-KBQA-U) is to drop the inappropriate repair step, and modify the LLM prompt to accommodate unanswerable questions, along with relevant in-context exemplars. Unlike in studies for unanswerability in general QA Slobodkin et al. (2023), we found that FuSIC-KBQA-U mostly generates incorrect logical forms for unanswerable questions.
As a remedy, we design a novel solution: FUn-FuSIC (Feedback for Unanswerability in FuSIC-KBQA). The key idea is to modify iterative repair, which earlier relied on a single strong verifier for the logical form’s incorrectness, to rely on a suite of strong and weak verifiers, where strong verifiers identify certain errors, whereas weak verifiers identify potential errors in the current logical form.
FUn-FuSIC’s verifiers consider both the logical form and the answer. For answers, non-emptiness check is now a weak verifier, given potentially unanswerable questions. For logical forms, we use strong verifiers to identify obvious syntactic and semantic errors. We also propose a novel verifier involving a 3-component LLM-based pipeline: non-equivalence of the original question and the back-translation of the logical form. This verifier is also weak, due to potential errors in back-translation as well as in equivalence classification. Using such iterative strong and weak verification based repair, FUn-FuSIC constructs a set of candidate logical forms. For selecting the consensus logical form from this set, using the majority answer as in self-consistency Wang et al. (2023) breaks down in the face of unanswerability. We introduce self-consistency for unanswerability, which assesses the likelihood of the majority answer, empty or otherwise, to select the consensus logical form.
Since no datasets exist for our novel task, we create two new datasets for KBQA transfer with unanswerable questions. Our experiments show that FUn-FuSIC comprehensively outperforms different categories of SoTA models suitably adapted for this task, including LLM-based and more traditional models. We further find that iterative repair of logical forms using weak verifiers holds promise for even for KBQA with only answerable questions. Using experiments over benchmark datasets for this task, we show that the restriction of FUn-FuSIC for the answerable setting improves upon the SoTA model for the task.
In summary, our specific contributions are as follows. (a) We propose the problem of few-shot transfer for KBQA with unanswerability. (b) We present FUn-FuSIC that uses iterative repair with error feedback from a diverse suite of strong and weak verifiers. (c) We create new datasets for the proposed task, which we make public. (d) We show that FUn-FuSIC outperforms adaptations of SoTA KBQA models for this new task. (e) We also show that even for answerable-only KBQA, FUn-FuSIC outperforms the corresponding SoTA model.
2 Related Work
In-domain KBQA using supervised models  Saxena et al. (2022); Zhang et al. (2022); Mitra et al. (2022); Wang et al. (2022); Das et al. (2022); Ye et al. (2022); Chen et al. (2021); Das et al. (2021); Shu et al. (2022); Gu et al. (2023) and using LLM few-shot approaches Li et al. (2023); Nie et al. (2024); Shu and Yu (2024) is well explored in literature. These use high volumes of labeled data, either for training or selecting the most relevant few shot exemplars.
For in-domain KBQA, unanswerability has recently been studied Patidar et al. (2023); Faldu et al. (2024).  Patidar et al. (2023) create the GrailQAbility dataset with different categories of unanswerability, and show the inadequacy of superficial adaptations of answerable-only KBQA models. RetinaQA Faldu et al. (2024) is the SoTA model for KBQA unanswerability. However, this also requires large volumes of training data.
For KBQA transfer Cao et al. (2022); Ravishankar et al. (2022), low-resource was originally not a focus. More recently, few-shot transfer for KBQA has been addressed by FuSIC-KBQA Patidar et al. (2024). FuSIC-KBQA uses a retrieve-then-generate framework with an LLM-based generation stage with iterative error feedback based repair. However, this formulation assumes answerability of all questions.
Simple LLM prompting techniques have been used to address unanswerability outside of KBQA Slobodkin et al. (2023), but without any notion of feedback or iterative repair. Other approaches  Shinn et al. (2023); Chen et al. (2023b) use execution based refinement for program generation but without any notion of unanswerability.
FUn’s iterative repair idea may be useful in other natural language to program generation tasks where non-existence of a program with the required specification has not been studied to the best of our knowledge, such as NL-to-SQL  Dong et al. (2023); Pourreza and Rafiei (2023) and program self-repair using LLMs Olausson et al. (2024); Madaan et al. (2023); Grattafiori et al. (2024).
3 Background & Problem Definition
Figure 1: Feedback with Unanswerability (FUn) and self consistency for Unanswerability (scUn) for a question when executed over different KBs with 4$\leq$ 4≤ 4 iterations. The question is answerable for KB3, but unanswerable for KB1 (schema incompleteness) and KB2 (data incompleteness). In the KB depictions, the top graph represents the schema with different node colors for different entity types, and the bottom graph represents the data. (Names of different (real) books related to the author entity in the question are abbreviated in the data graph.) FUn iterations are shown in the gray blocks, with Strong Verifiers named using red text and Weak Verifiers using green text and $i$ denotes the iteration number. The outcome of verification is denoted as True or False, and the non-empty answer is shown when V4b returns False. The Syntax Error Verifier (V1) is omitted for brevity. scUn is shown in the blue blocks. The candidate logical forms ($L$) for sCun are shown at the top. For Non-empty Ans and Empty Ans agreement checks for, the outcome is either ‘Consensus’ or ‘No consensus’. At the end, FUn-FuSIC ‘Return’s a logical form (possibly NK) and an answer (possibly NA).
A Knowledge Base (KB) $G$ consists of a schema and data. The schema consists of entity types (or classes) $T$ and binary relations $R$ defined over pairs of types. The data consists of entities $E$ as instances of types $T$, and triples or facts $F\subseteq E\times R\times E$. Given a target KB $G^{t}$ and a natural language question $q^{t}$, the basic KBQA task is to generate a structured query or logical form $l^{t}$ (in a KB query language, such as SPARQL), which when executed over $G^{t}$ returns an answer $A^{t}$ ($A^{t}\subset E$ in general). Other than SPARQL  Patidar et al. (2024), niche languages such as s-expressions  Li et al. (2023); Gu et al. (2023) are commonly used for logical forms in KBQA. In supervised in-domain KBQA, the target has large volumes of labeled training examples of questions and associated logical forms. For few-shot in-domain KBQA in contrast, target few-shots $D^{t}$ contain tens of labeled training examples. In few-shot transfer learning for KBQA Patidar et al. (2024), a related source domain has a source KB $G^{s}$ (with its own types, relations, entities and facts), and a larger source training set $D^{s}$ with thousands of labeled training examples.
Following Patidar et al. (2023), a question $q$ is answerable for a KB $G$ if it admits a corresponding logical form $l$ which when executed over $G$ returns the ideal non-empty answer $A$. A question is unanswerable if it either (a) does not have a valid logical form for $G$ (schema-level unanswerability), or (b) it has a valid logical form $l$ for $G$, but $l$ returns an empty answer upon execution on $G$, different from the ideal non-empty answer (data-level unanswerability). Schema-level unanswerability arises due to missing types and relations, while missing entities and facts lead to data-level unanswerability. However, absence in $G$ of any entity mentioned in the question is categorized as schema level unanswerability, since it invalidates the logical form. More details are in Appendix  A.1.2.
In KBQA with unanswerability Patidar et al. (2023); Faldu et al. (2024), given a question $q$, the model needs to output (a) a logical form $l$ and a non-empty answer A for answerable $q$, (b) $l=$ NK (No Knowledge) for schema-level unanswerable $q$, or (c) a valid logical form $l$ and $a=$ NA (No Answer) for data-level unanswerable $q$. In the supervised in-domain setting, this task involves large volumes of labeled training questions, containing both answerable and unanswerable, for the target.
We now define our problem of interest: few-shot transfer learning for KBQA with unanswerability. A target question $q^{t}$ may be answerable or unanswerable due to missing schema or data in the target KB $G^{t}$. Target few-shot examples $D^{t}$ contain both answerable and unanswerable questions of different categories. The source training data $D^{s}$ has large volumes of labeled training data. Considering real world constraints, where most KBQA datasets contain only answerable questions, we assume that $D^{s}$ contains only answerable questions. Compared to the earlier few-shot KBQA transfer task definition, now there is additionally an unanswerability mismatch between the source and the target distribution. More details are in the Appendix (Sec.  A.1).
4 Proposed Approach: FUn-FuSIC
Our proposed model FUn-FuSIC preserves the basic architecture of FuSIC-KBQA and adapts its iterative repair strategy for unanswerability. The high-level algorithm is described in Algo. 1. (Since the algorithms are not specific to the transfer task, we use $q$ instead of $q^{t}$ for brevity.) Preserving the retrieve-then-generate framework of FuSIC-KBQA, the retrieval stage (line 2) performs KB retrieval for $q^{t}$ using a set $R$ of one or more supervised retrievers. Each retriever $R_{i}$ is source-trained and further target fine-tuned if required. The retrieval output $r$ of each $R_{i}$ consists of relevant schema elements (types and relations) for $q^{t}$, and data paths emanating from mentioned entities in $q^{t}$. The union of these, along with $q^{t}$, is fed to the generation stage, which uses prompting with an LLM $\mathcal{L}$ to generate logical forms using the target few-shots $D^{t}$. More details of the retrieval stage are in the Appendix (Sec. A.8.1).
[…]
6:else return $\mbox{scUn}(q,L,\mathcal{L})$
FUn-FuSIC differs from FuSIC-KBQA in its iterative repair strategy in lines 3, 4 and 6 of Algo. 1. The LLM generation instruction is modified to admit the possibility of unanswerability, and the few shots are modified to include examples of unanswerable questions. However, this simple approach is error-prone. So, we bias the instruction towards one type of error. Specifically, when uncertain about answerability of the question, Prompting for Unanswerability (PUn) (line 3) instructs $\mathcal{L}$ to generate a (possibly incorrect) logical form instead of $l=\texttt{NK}$. The detailed prompt $I$ is in the Appendix (Sec. A.11.1).
We now come to the more significant modifications. First, $l^{(0)}$ is iteratively repaired using feedback as before, but this step is adapted for unanswerability. This iterative repair, which we name Feedback for Unanswerability (FUn) (line 4), either confidently outputs a single logical form $l$ (with corresponding answer $A$) (line 5) or generates a set $L$ of candidate logical forms, which is further analyzed for a consensus logical form and answer. For this, we introduce self-consistency for Unanswerability (scUn) (line 6). scUn assesses the likelihood of the majority answer in $L$, empty or otherwise, to produce the final output. In the rest of this section, we describe FUn and then scUn. Fig. 1 illustrates flow of FUn and scUn using examples. A real example of FUn execution is in Sec. A.12.
Algorithm 2 $\mbox{FUn}(q,l,n,V^{s},V^{w},G,\mathcal{L})$
[…]
2:while $i\leq n$ do
3:  for $j=1\mbox{ to }k_{1}$ do
4:   $(e,f)=V^{s}_{j}(l^{(i)},q,G)$
5:   $F=\mbox{Append}(F,f)$
6:   $\mbox{{\bf if} }(!e)\mbox{ {\bf break}}$
7:  end for
8:  for $j=1\mbox{ to }k_{2}$ do
9:   $(e,f)=V^{w}_{j}(l^{(i)},q,G)$
10:   $F=\mbox{Append}(F,f)$
11:   $\mbox{{\bf if} }(e)\mbox{ }L=L\bigcup\{l^{(i)}\}$; $k++$
12:  end for
13:  $i=i+1$
14:  $l^{(i)}=Gen(\mathcal{L},I,q,F)$
[…]
16:end while
[…]
The FUn algorithm is described in Algo. 2 Starting with the initial logical form $l^{(0)}$, FUn performs at most $n$ verify-and-repair iterations to create a candidate set $L$ of probable logical forms. Fig. 1 shows 3 FUn iterations for KB1 and KB2, and 2 for KB3. In the $i^{th}$ iteration, FUn generates a new logical form $l^{(i)}$ by prompting $\mathcal{L}$ using $q^{t}$ and feedback $F$ received from checks in all previous iterations (line 13). $l^{(i)}$ goes through a sequence of verifications. FUn uses two sets of verifiers. The strong verifiers $V^{s}$ are guaranteed to be correct, while the weak verifiers $V^{w}$ are potentially erroneous. $k_{1}$ and $k_{2}$ denote the total number of strong and weak verifiers respectively. The specific verifiers that we use in this paper are defined later in the section. A template-based feedback string $f$ is appended to the generation prompt for $l^{(i+1)}$ based on the specific verifier that $l^{(i)}$ failed. If $l^{(i)}$ fails a strong verifier, it is rejected (line 6). In the example, this happens for all three KBs in iteration 1. If $l^{(i)}$ passes all checks, strong and weak (line 15), FUn terminates by outputting $(l=l^{(i)},A=A^{(i)})$, where $A^{(i)}$ is the answer obtained by executing $l^{(i)}$ (denoted $\mbox{X}(l^{(i)},G)$). This happens in iteration 3 for KB3. Otherwise, if $l^{(i)}$ passes at least one weak verifier but not all, it is added to candidate logical form set $L$ (line 11). This happens for iterations 2, 3 and 4 for KB1 and KB2.
Logical form verifiers:
FUn uses a suite of verifiers, categorized as strong ($V^{s}$) and weak ($V^{w}$). These may be syntactic, semantic, or execution-based, defined using simple rules or complex LLM functions over $l$, $q$ and $G$. Note that unlike unit tests in program synthesis, the verifiers do not have knowledge of the gold logical form, the gold answer or answerability of the question.
We now briefly describe the specific verifiers that we use for this paper. Additional details about the verifiers are in the Appendix (Sec. A.11.2 and Sec. A.3). Note that FUn is a framework that is capable of working with a wholly different suite of meaningful verifiers.
(V1) Syntax Error:
As in FuSIC-KBQA, this verifier executes the logical form $l$ over $G$ and checks for syntax error. This is a strong check — a valid logical form cannot have syntax error.
(V2) KB Inconsistency:
A logical form $l$ may be inconsistent with the schema of $G$. We identify semantic errors of different categories, such as type-incompatibility and schema hallucinations, implemented using rules over $l$ and $G$. These are also strong verifiers.
(V3) Question-Logical Form Disagreement:
This verifier checks if $l$ is semantically equivalent to the original natural language question $q$. In Fig. 1, LF3 for KB1 disagrees with $q$. This is a weak verifier. First, $q$ may not have any equivalent logical form for $G$ due to intrinsic ambiguities even when it is answerable. For example, $q$ mentions a person from a country, when $G$ has the relations born in and works in between these types. Secondly, this verifier is a probabilistic classifier that naturally makes occasional mistakes. We define equivalence check between $l$ and $q$ using a multi-stage LLM pipeline, involving naturalization of $l$ to $l^{n}$, back-translation of $l^{n}$ to natural language question $q^{b}$ and semantic equivalence check between $q$ and $q^{b}$. More details are in the Appendix (Sec.A.3).
(V4) Answer Inconsistency:
This verifier executes $l$ over $G$ to obtain an answer $A$ and then checks its compatibility with $q$. This may fail for different reasons, such as (V4a) $A$ containing an entity mentioned in $q$, (V4b) $A$ being empty, and others. Note that V4a is a strong verifier while V4b is weak, since an empty answer is valid for unanswerable questions (as for LF3 for KB2), but invalid for answerable ones.
Identifying Candidate Logical Forms:
Unless some logical form passes all checks and is therefore returned (Algo. 2 line 15), FUn constructs a candidate set $L$ of logical forms that are potentially flawed but not certainly so. For our specific suite of weak verifiers, $l^{(i)}$ is added to $L$ if it passes one of V4b ($A$ is non-empty) as for LF3 for KB1, or V3 ($l^{i}$ is equivalent to $q$), as for LF3 for KB2.
Self Consistency for Unanswerability (scUn):
Given a candidate set $L$ of logical forms and a question $q$, scUn assesses if the best candidate $l^{*}\in L$ has sufficient confidence. If so, it outputs ($l=l^{*}$, $A=A^{*}$), as for KB2, $A^{*}$ being the answer from executing $l^{*}$ (may be NA). Otherwise, scUn outputs ($l=$ NK, $A=$ NA), as for KB1. For identifying the consensus choice from $L$, one possibility is self-consistency (sc) Wang et al. (2023); Chen et al. (2023a) that considers the answer for each $l\in L$, and returns those with the most common answer. This requires some answer to accumulate enough probability by aggregation over reasoning paths. However, for unanswerable questions, no single answer accumulates sufficient probability, and sc returns some low probability answer.
To address this, scUn first identifies via execution the most popular non-empty answer $A^{*}$ among logical forms in $L$, and decides using a threshold $t$ if it has enough supporters in $L$ (we use $t=\lfloor\frac{|L|}{2}\rfloor$). If so, scUn uses LLM prompting to select the most appropriate supporting logical form $l^{*}\in L$ considering $q$, and outputs ($l=l^{*}$, $A=A^{*}$). However, for KB1, the 3 logical forms among the candidates have 3 different answers, and therefore no consensus emerges ($\lfloor\frac{|L|}{2}\rfloor=1$). Here, scUn considers logical forms from $L$ that agree on $A=$ NA. If there are multiple such candidates, scUn selects the most suitable candidate $l^{*}$, again using LLM prompting, and outputs ($l=l^{*}$, $A=$ NA). If there is no such candidate, scUn outputs ($l=$ NK, $A=$ NA). For KB2 in the example, scUn selects LF3 – the only logical form with empty answer. Further details on scUn are in the Appendix (Sec. A.2).
5 Experiments
WebQSP $\rightarrow$ GrailQAbility WebQSP $\rightarrow$ GraphQAbility
Model Overall Answerable Unanswerable Overall Answerable Unanswerable
F1 EM-s F1 EM-s F1(L) F1(R) EM-s F1 EM-s F1 EM-s F1(L) F1(R) EM-s
RetinaQA 58.4 42.2 28.7 26.0 88.0 84.8 58.4 49.7 35.8 18.7 15.2 80.7 78.7 56.4
Pangu 54.5 43.8 31.2 29.6 83.8 80.4 58.0 53.4 33.0 30.3 26.4 76.5 74.8 39.6
FuSIC-KBQA-U 76.6 48.2 67.5 59.2 85.6 80.4 37.2 67.5 34.8 49.3 40.0 85.7 82.8 29.6
KB-Binder 43.7 33.0 19.5 16.5 67.9 66.5 49.5 44.3 36.1 27.5 21.6 61.0 61.0 50.7
FUn-FuSIC 76.6 60.2 67.1 61.2 85.1 80.0 59.2 70.0 53.8 50.7 42.8 89.2 86.5 64.8
Table 1: Performance of different models on two datasets for few-shot KBQA transfer with unanswerability. Answerable and Unanswerable record performance for corresponding subsets and Overall for the entire dataset.
We now present experimental evaluation of FUn-FuSIC. First, for few-shot KBQA transfer with unanswerability, we address the following research questions. (R1) How does FUn-FuSIC compare against SoTA KBQA models suitably adapted for this setting? (R2) How does FUn-FuSIC perform across different categories of unanswerability? (R3) How do the different components of FUn-FuSIC contribute to its performance? Then, for answerable KBQA few-shot transfer, we ask: (R4) How does FUn-FuSIC compare against SoTA KBQA models for this setting?
5.1 Experimental Setup
Datasets:
For in-domain and answerable KBQA, the three most popular datasets are GrailQA Gu et al. (2021), GraphQA Su et al. (2016) and WebQSP Yih et al. (2016). All of these have the same back-end KB (Freebase). For few-shot KBQA transfer, the only available datasets also have only answerable questions Patidar et al. (2024). GrailQAbility is the only available KBQA dataset with unanswerable questions Patidar et al. (2023). This was constructed starting from GrailQA Gu et al. (2021) by systematically deleting schema and data elements from the back-end KB to introduce different categories of unanswerability into the queries.
[…]
WebQSP $\rightarrow$ GrailQAbility WebQSP $\rightarrow$ GraphQAbility
Model Schema Level Data Level Schema Level Data Level
F1(L) F1(R) EM-s F1(L) F1(R) EM-s F1(L) F1(R) EM-s F1(L) F1(R) EM-s
RetinaQA 94.1 90.9 79.4 76.3 72.9 14.1 83.2 82.0 72.3 73.7 72.7 12.1
Pangu 91.1 87.9 87.9 69.6 65.9 00.0 77.3 74.4 74.4 73.3 72.7 00.0
FuSIC-U 85.4 80.6 30.9 86.0 80.0 49.4 86.6 82.6 19.0 83.3 83.3 51.5
KB-Binder 75.1 73.9 70.1 53.1 51.5 09.5 67.0 65.9 60.9 41.2 41.2 06.8
FUn-FuSIC 85.8 81.2 70.9 83.8 77.6 36.5 92.4 87.5 75.6 80.3 80.3 34.8
Table 2: Model performance for categories of unanswerable questions. FuSIC-U is short hand for FuSIC-KBQA-U.
Models for comparison:
As few-shot transfer for KBQA with unanswerability is a novel task, there are no existing baselines. For in-domain KBQA with unanswerability, RetinaQA Faldu et al. (2024) and the unanswerability-adapted version of Pangu Gu et al. (2023) are the SoTA models. For these, we use the available code. [^2] https://github.com/dair-iitd/RetinaQA, [^3] https://github.com/dki-lab/Pangu More details are in the Appendix (Sec. A.8.3).
FuSIC-KBQA is the SoTA model for few-transfer for KBQA with only answerable questions. KB-Binder Li et al. (2023) is the SoTA for in-domain few-shot KBQA. Overall, FuSIC-KBQA and KB-Binder outperform all other supervised and LLM-equipped KBQA models adapted for few-shot transfer Patidar et al. (2024). We use available code for KB-Binder [^4] https://github.com/ltl3A87/KB-BINDER, and our own implementation for FuSIC-KBQA. To adapt these two baselines for unanswerability, for fair comparison, we modify their logical form generation prompt in the same fashion as PUn for FUn-FuSIC. Additionally, for FuSIC-KBQA, we remove execution-guided feedback (EGF) since it fails for unanswerability. We denote this model FuSIC-KBQA-U. Observe that FuSIC-KBQA-U can also be seen as an ablation of FUn-FuSIC, without FUn. More details about KB-Binder and FuSIC-KBQA are in the Appendix (Sec. A.9.1).
We use $\mathcal{L}=$gpt-4-0613 for all LLM-equipped models. For fair comparison, we allocate to all such models the same maximum aggregated prompt length for a question. This is satisfied by equipping FUn-FuSIC with zero-shot generation and $n=4$ FUn iterations, FuSIC-KBQA-U with 5-shot generation and KB-Binder with 25-shot generation.
Though FUn-FuSIC and FuSIC-KBQA allow flexible use of multiple supervised retrievers, for meaningful comparison with RetinaQA, we adapt RetinaQA as retriever for FUn-FuSIC and FuSIC-KBQA-U. More details about FuSIC-KBQA’s retriever and compute infrastructure are in the Appendix (Sec. A.8.2).
Evaluation Measures:
For KBQA as semantic parsing task, evaluation of logical forms is primary. For this, the existing EM measure Ye et al. (2022) is defined only for logical forms represented using s-expressions. FuSIC-KBQA-U and FUn-FuSIC output logical forms in SPARQL, and Pangu, RetinaQA and KB-Binder in s-expression. So we propose a new measure EM-s that checks approximate equivalence for a pair of programs either in SPARQL or s-expression. More details are in the Appendix (Sec. A.5).
As in standard KBQA evaluation, we also evaluate answers. This is a secondary evaluation for giving the benefit of the doubt for getting the right answer, possibly via a logical form not equivalent to the gold-standard according to EM-s. For answer evaluation in KBQA with unanswerability,  Patidar et al. (2023) introduced lenient F1, denoted F1(L), in addition to regular F1, denoted F1(R). F1(L) relaxes F1(R) by not penalize the original answer for the complete KB. Note that obtaining the right answer by chance has much higher probability than for logical forms, particularly for unanswerable questions with NA as the correct answer.
WebQSP $\rightarrow$ GrailQAbility WebQSP $\rightarrow$ GraphQAbility
Model Answerable Schema L. UnAnswerable Data L. UnAnswerable Answerable Schema Level UnAns Data Level UnAns
F1 EM-s F1(L) F1(R) EM-s F1(L) F1(R) EM-s F1(L) EM-s F1(L) F1(R) EM-s F1(L) F1(R) EM-s
FUn-FuSIC 74.0 70.0 90.9 87.9 75.8 64.7 64.7 11.8 59.0 48.0 97.1 91.4 80.0 86.7 86.7 26.3
scUn $\Rightarrow$ sc 74.0 70.0 73.0 72.7 33.3 58.8 52.9 23.5 64.0 54.0 74.4 68.6 22.9 80.0 80.0 13.3
w/o syntax 67.3 64.0 87.9 90.9 42.4 76.5 70.6 17.7 57.0 46.0 81.0 76.9 26.9 75.8 75.0 16.7
w/o kb-inc 70.3 66.0 90.9 90.9 9.1 76.5 70.6 29.4 51.7 42.0 100.0 100.0 4.2 79.5 75.0 20.8
w/o q-lf 71.0 68.0 69.7 63.6 0.0 41.2 35.3 5.9 47.7 38.0 69.5 65.4 07.7 67.0 62.5 20.8
w/o ans-inc 72.0 70.0 85.9 84.9 33.3 70.6 70.6 23.5 55.0 44.0 80.8 76.9 26.9 79.5 79.2 25.0
Table 3: Ablation performance of FUn-FuSIC (removing individual components with replacement) on subset of WebQSP $\rightarrow$ GraphQAbility. scUn $\Rightarrow$ sc denotes replacing scUn with self consistency. Other rows remove verifiers for syntax error (w/o syntax) (V1), KB inconsistency (w/o kb-inc) (V2), question logical form disagreement (w/o q-lf) (V3) and answer incompatibility (w/o ans-inc) (V4). Evaluations are on 100 instances from test sets (50 answerable and 50 unanswerable questions sampled uniformly at random)
5.2 Unanswerability Setting
We first address research question R1. Performances of different models for few-shot transfer with unanswerability are recorded in Tab. 1. First, we observe that FUn-FuSIC significantly outperforms all baselines in terms of EM-s, and performs at par with FuSIC-KBQA-U and significantly better than all other models in F1. The other LLM-equipped models are not significantly better than the supervised models. All the baselines perform almost at par for GraphQAbility, and KB-Binder performs worse than the other 3 for GrailQAbility. This establishes usefulness of FUn equipped with scUn for few-shot transfer KBQA with unanswerability, beyond LLM-usage. Secondly, each model trades off performance differently between answerable and unanswerable questions. RetinaQA, Pangu and also KB-Binder fare better for unanswerable questions, while FUn-FuSIC and FuSIC-KBQA-U fare better for answerable ones. However, FUn-FuSIC achieves the best balance across the two subsets.
We next briefly address research question R2. Performance of different models for different categories of unanswerability are recorded in Tab. 2. All models struggle to fare well simultaneously for data-level and schema-level unanswerability. FuSIC-KBQA-U performs the best for data-level while performing poorly (in terms of EM-s) for schema-level. Conversely, RetinaQA performs well for schema-level, but has poor data-level EM-s. FUn-FuSIC outperforms other models in schema-level unanswerability while being slightly worse in data-level unanswerability. But among all models, it achieves the best tradeoff by far across unanswerability categories.
We next address research question R3. Tab. 3 records the ablation analysis of FUn-FuSIC. We see that the KB-Inconsistency verifier (V2) and the Q-LF Disagreement verifier (V3) lead to significant improvements in EM-s. The biggest benefit comes from V3 for both answerable and unanswerable questions. Without V3, the correct LF for schema-level unanswerability is almost never generated, though answer accuracy stays high, indicating inability to reason. Similarly, removing V2 reduces EM-s for schema-level unanswerable questions, with answer accuracy remaining high, indicating that answers are often correct despite flawed logical forms. The Answer Incompatibility verifiers (V4) also makes significant contributions to the performance. This analysis highlights the necessity of a mix of weak and strong verifiers for structural and semantic validity. Beyond verifiers, replacing scUn with self-consistency, as expected, leads to a drastic drop in unanswerable performance (though this comes with a benefit for answerable questions).
WebQSP $\rightarrow$ WebQSP $\rightarrow$
Model GrailQA-Tech GraphQA-Pop
FuSIC-KBQA 70.8 52.3
FUn-FuSIC(sc) 73.6 67.0
FuSIC-KBQA-U 62.6 43.4
FUn-FuSIC(scUn) 71.2 65.0
Table 4: Performance using F1 of different models for few-shot KBQA transfer with only answerable questions. The models in the top block have prior knowledge of answerability, while those in the bottom block do not.
[…]
5.3 Answerable Setting
We now address research question R4 for answerable-only KBQA transfer. We use two datasets from existing literature Patidar et al. (2024), including the hardest one (WebQSP $\rightarrow$ GraphQA-Pop). [^5] https://github.com/dair-iitd/FuSIC-KBQA/ For enabling comparison with earlier results, we use TIARA Shu et al. (2022) as the retriever for all models in this experiment.
This setting admits two sub-cases: (A) the models have knowledge that all questions are answerable, and (B) though all questions are answerable, the models do not have this knowledge.
Setting (A) has been studied for KBQA Patidar et al. (2024), and FuSIC-KBQA is the established SoTA model, outperforming a host of supervised and LLM-based models adapted for the task. To adapt for this setting, FUn-FuSIC requires three simplifications. (i) PUn is replaced with prompt for answerability, (ii) In FUn, V4b (empty answer) is moved from the set of weak verifiers to that of strong verifier, and (iii) scUn is replaced by standard self-consistency.
The first two rows in Tab. 4 record performance for setting (A). FUn-FuSIC significantly outperforms FuSIC-KBQA on both datasets, creating a new SoTA for this setting. This shows the usefulness of iterative repair with a suite of strong and weak verifiers followed by self-consistency for few-shot KBQA transfer, even without unanswerability.
In the more realistic setting (B), which has not been studied before, the models make predictions assuming unanswerability. Here, we evaluate FuSIC-KBQA-U and FUn-FuSIC as in Sec. 5.2, only there are no truly unanswerable questions. The bottom two rows of Tab. 4 record the performance of the two models in this setting. We see that FUn-FuSIC outperforms FuSIC-KBQA by a very large margin. This further establishes the usefulness of scUn when guarantees about answerability are not available.
5.4 Error Analysis
EM-s <1<1< 1 46.2
Retr. Err. 23.4
Gen. Err. 22.8
$l^{*}=$ NK, $^{\hat{l}\neq}$ NK 8.4
$l^{*}\neq$ NK, ${\hat{l}=}$ NK 4.6
$l^{*}\neq$ NK, $^{\hat{l}}\neq$NK, $l^{*}\neq{\hat{l}}$ 9.8
Table 5: FUn error analysis on WebQSP $\rightarrow$ GraphQAbility. $l^{*}$ $^{\hat{l}}$ denote gold & generated logical forms. Retrieval error means retrieval $r$ is missing 1$\geq$ 1≥ 1 KB elements (class, relation, entity) necessary for $l^{*}$. Generation error implies ${\hat{l}}\neq l^{*}$ despite correct retrieval.
For WebQSP $\rightarrow$ GraphQAbility, we analyzed questions for which logical forms generated by FUn-FuSIC are incorrect (EM-s <1<1< 1). Results are in Tab. 5. We found three main causes for generation errors. (1) Some questions are inherently ambiguous, admitting multiple valid logical forms $l_{1}$ and $l_{2}$ in the original complete KB, though only one is recognized as the gold ($l^{*}=l_{1}$). Deletion to introduce unanswerability eliminates $l_{1}$, so that that $l^{*}=$ NK, and the prediction ${\hat{l}}=l_{2}$ is unfairly penalized. (2) Here, $l^{*}=l_{1}$ and the prediction ${\hat{l}}=l_{2}$, such that $l_{1}\neq l_{2}$ but are semantically equivalent. $l_{1}$ and $l_{2}$ are incorrectly judged non-equivalent by EM-s. (3) Here, FUn is unable to generate $l^{*}$ or any semantic equivalent of it within its iteration limit.
6 Conclusions
For real-world robust and low-resource KBQA, we have proposed the novel task of few-shot transfer learning with unanswerability. We have introduced a new notion (FUn) of iterative feedback guided repair for answerable as well as unanswerable questions. FUn (i) uses feedback from a diverse suite a strong and weak verifiers – including a novel back-translation based verifier – to create a set of candidate logical forms, and (ii) assesses this candidate set to either to detect unanswerability (and its category) or identify the best logical form using self consistency adapted for unanswerability (scUn). We propose FUn-FuSIC that replaces the existing the iterative strategy, that assumes answerability of questions, with FUn in the SoTA few-shot answerable-only KBQA transfer model (FuSIC-KBQA). Using two newly created datasets for this novel task, we show that FUn-FuSIC significantly outperforms adaptations of FuSIC-KBQA and other SoTA models for this setting, and also for answerable few-shot transfer KBQA.
Our error analysis suggests that performing well across categories of unanswerability for few-shot transfer is still a challenge for KBQA and should be a focus of further research. We have made our datasets and other resources public  [^6] https://github.com/dair-iitd/FUn-FuSIC.
Limitations
Since LLM inference involves randomness, experiments should ideally be repeated for multiple runs and results should report averages and error bars. Unfortunately, we were not able to do this due to the prohibitive cost of GPT-4, and our results are based on single runs.
While GPT-4 is currently the best performing LLM, it is proprietary as well as expensive. Ideally, evaluation should include open-source freely accessible LLMs as well. We expect performance of all LLM-based approaches to drop when GPT-4 is replaced by a less powerful, open LLM. Nonetheless, earlier research has shown that models with Mistral instead of GPT-4 still outperform fully supervised models for answerable few-shot transfer Patidar et al. (2024). Whether this trend holds for the unanswerable setting is an open question. That said, following current trends, we expect the ability of open LLMs to steadily improve in the coming years.
Risks
At the highest level, our work reduces risk compared to existing KBQA systems, which when inadequately adapted in a low-resource setting, incorrectly answer unanswerable questions, without acknowledging lack of knowledge. However, can incorrectly inferring unanswerability, citing lack of knowledge when knowledge is in fact available, be a new type of risk? While we cannot imagine such a risk at the present time, this may require more careful consideration. In any case, KBQA models for unanswerability should strive to minimize this type of error, along with the other types.
Acknowledgments
Mausam is supported by a contract with TCS, grants from IBM, Verisk, Huawei, Wipro, and the Jai Gupta chair fellowship by IIT Delhi. Indrajit would like to thank KnowDis AI for supporting his participation in the conference. Riya would like to thank Graviton Research Capital for supporting her participation in the conference. The authors would like to thank the IIT-D HPC facility for its computational resources. We also thank Microsoft Accelerate Foundation Models Research (AFMR) program that provided us access to OpenAI models. We are also thankful to Mayur Patidar for helpful discussions.
Appendix A Appendix
A.1 KBQA Elaboration
Here we elaborate on different aspects of the KBQA problem.
A.1.1 Challenges in Few Shot Transfer Learning for KBQA
The source and target tasks may differ significantly. First, the data and schema of the knowledge bases $G^{t}$ and $G^{s}$ and the domains they cover may be different. Secondly, the distributions of questions and logical forms defined over the KBs may be different in $D^{t}$ and $D^{s}$.
A.1.2 Different Types of Unanswerability
Unanswerable questions in KBQA can be categorized into (a) Schema Level Unanswerability : the question does not have a corresponding logical form that is valid for the KB , (b) Data level unanswerability: it has a valid logical form $l$ for the KB, but which on executing returns an empty answer. Schema level unanswerable questions can further be categorized into (1) Missing Class: The class/type required to construct the logical form is not defined for the KB, (2) Missing Relation: The relation required to construct the logical form is not defined for the KB, (3) Missing Topic Entity: The topic entity specified in the question is missing from the KB. Data level unanswerable questions can be categorized into (1) Missing entity: all classes and relations required to construct the logical form are present in the KB , but there exists no path from the topic entity node to the answer node in the KB due to missing intermediary entities (2) Missing Fact: all classes, relations and entities required to answer the question are present in the KB. However, the (subject, relation, object) path is not connected in the KB.
A.2 Algorithm for Self Consistency with Unanswerability
The high level algorithm for self consistency with Unanswerability (scUn) is described in Algo. 3.
[…]
The high level algorithm for assessing confidence in the set of candidate logical forms (assessConf) is described in Algo. 4. We use abbreviations NE and E to indicate non-empty and empty respectively. popAnsNE is abbreviation of ”(most) popular answer non-empty”.
Algorithm 4 $\mbox{assessConf}(q,L,\mathcal{L})$
[…]
2:if $(c)$ then
3:  $l=\mbox{selectBestNE}(q,L^{p},\mathcal{L})$
4:  $\mbox{{\bf return}}(\mbox{T},l,A^{p})$
5:end if
[…]
7:if $(c)$ then
8:  $l=\mbox{selectBestE}(q,L^{p},\mathcal{L})$
9:  $\mbox{{\bf return}}(\mbox{T},l,\texttt{NA})$
10:end if
[…]
A.3 Details of FUn Verifiers
Here we discuss the verifiers V2, V3 and V4 in more detail.
(V2) Semantic Error (KB Inconsistency):
A syntactically correct logical form $l$ may still be inconsistent with the schema of $G$. This is a likely error even for SoTA LLMs since these are unfamiliar with the specific KB $G$. Semantic errors have different categories, such as type-incompatibility, schema hallucinations. (V2a) Incompatibility in types: $l$ contains a variable and a connecting relation whose types are incompatible in $G$. This is the case for LF1 for all three KBs in the example. (V2b) Schema hallucinations: $l$ contains schema elements (types, relations, entities) absent in $G$. (V2c) Type casting errors: Literals in $l$ are not correctly type cast for $G$, e.g. numeric literals as float for Freebase. All of these are certain checks, and are implemented using rules defined over $l$ and $G$. The feedback mentions the type of error and the specifics, e.g., the hallucinated relation, or the incompatible type-relation pair.
(V3) Question-Logical form Disagreement:
FUn performs equivalence check between $l$ and $q$ using a novel multi-stage LLM pipeline. (i) The variable names in $l$ are first naturalized to $l^{n}$ considering $q$ and preserving semantics, e.g. by replacing ’?x’ with ’?actor’. (ii) $l^{n}$ is back-translated into a natural language question $q^{b}$. (iii) $q^{b}$ is finally checked for semantic equivalence with $q$. The first two steps are performed using zero-shot prompting, while the last is performed using few-shots constructed using the target few-shots $D^{t}$. The feedback mentions lack of equivalence as the type of error.
(V4) Answer Inconsistency:
If the $l$ is syntactically and semantically correct, it is executed over $G$ to obtain an answer $a$. $a$ is then checked for compatibility with $q$. This may fail for different reasons. (V4a) $a$ (which is a set in general) contains an entity also in $l$ and therefore mentioned $q$, which is an aberration. (V4b) $a$ is empty, as in LF3 for KB2 in the example. All of these checks are implemented using rules defined over $l$ and $G$. Note that while the first two are certain checks, the last is not. An empty answer is valid for unanswerable questions, as for LF3 for KB2, but invalid for answerable ones. As before, the feedback mentions the type of error and the specifics.
Knowledge Base Logical Forms NL Questions
[…]
SrcAv/TgtAv Cosine Sim SrcAv/TgtAv
[…]
Table 6: Statistics for different source and target (test set) KBQA task pairs in terms of the knowledge base, logical forms and natural language questions. ‘Domain(JS)’ is Jensen Shannon (JS) divergence between domain distribution of questions, ‘New Rel%’ is percentage of questions in target with new (unseen) relations, ‘Function (JS)’ is JS-divergence between distributions over functions in logical forms, ‘#\##Relations’ shows the source average and the target average for number of relations per logical form, ‘Src-Tgt Cosine Sim’ is average minimum cosine distance between source and target questions and ‘#\##Tokens’ shows the source average and target average of number of tokens per question.
A.4 Additional Dataset Statistics
[…]
Dataset Schema level Data level
UnAns UnAns
[…]
Table 7: Percentages of schema and data level unanswerability among unanswerable questions.
A.5 EM-s: Automated Approximate Equivalence Check for SPARQL
As has been observed in Patidar et al. (2023), answer evaluation by itself is not a robust measure for evaluation of KBQA models when the dataset contains unanswerability. Traditional KBQA models that generate s-expressions can be evaluated using EM, which checks for logical form equivalence between two logical forms, since it is possible to compare equivalence between two s-expressions efficiently. However, FUn-FuSIC generates SPARQL queries instead. Directly comparing program equivalence between two SPARQL queries is an undecidable problem  [^7] https://users.dcc.uchile.cl/~cgutierr/papers/expPowSPARQL.pdf. Patidar et al. (2024) suggest a semi-automatic strategy for comparison of sparql queries. We propose a completely automatic metric for SPARQL equivalence check. Two SPARQL queries are equivalent by the EM-s check if (a) the relations occurring in the two queries are same. (b) the entities occurring in the two queries are the same (c) the answer set obtained by executing the queries over the KB are the same. Note that the EM-s check is necessary, but not sufficient for two SPARQL queries to be equivalent.
[…]
A.6 Model Evaluation: Additional Details
In this section, we do a deeper evaluation of performance of different models across different categories of unanswerability, as explained in Patidar et al. (2023). There are two broad categories of unanswerability — schema level unanswerability (absence of knowledge in terms of KB ontology or entities required to construct the logical form) and data level unanswerability (absence of facts or intermediate entities of the logical form path on the KB).
We expect that (a) due to poor ability of supervised models to generalize in transfer learning settings, RetinaQA will struggle to generate correct logical forms for data level unanswerable questions, and (b) due to the strong generalization ability of FuSIC-KBQA, it should be able to perform well for data level unanswerable questions. However, since it is biased towards returning incorrect logical forms instead of abstaining from returning a logical form, it will perform poorly at identifying schema level unanswerable questions. (c) FUn-FuSIC should be able to maintain the performance of FuSIC-KBQA on data level unanswerable questions to a large extent, while significantly improving the performance on schema level unanswerable questions.
[…]
A.7 FUn Cost Analysis for Proprietary LLMs
FuSIC-KBQA, as well as the adapted versions of FuSIC-KBQA, such as FuSIC-KBQA-U and FUn-FuSIC rerank the classes, relations and paths. The total cost for reranking for one question is \$0.16. The cost for generation of logical form from a prompt with 5 in-context examples is \$0.16. Thus, the approximate cost for inference of one question by FuSIC-KBQA-U is \$0.32.
The cost for generation of logical form from a prompt with 0 in-context examples is \$0.04. The cost of checking whether two natural language questions are equivalent or not, using few-shot exemplars and chain of thought prompting is also \$0.04. The approximate cost of inference of one question by FUn-FuSIC varies between \$0.24 and \$0.48. The average cost over 50 randomly sampled questions from the test set is around \$0.34.
Hence, the two models are comparable in terms of cost.
A.8 Model Adaptation Details
Here we discuss adaptation details for the models that we have built upon (FuSIC-KBQA), used as retrievers (RetinaQA) and for comparison.
A.8.1 FuSIC-KBQA Details
Our proposed approach FUn-FuSIC builds upon the the base architecture of FuSIC-KBQA Patidar et al. (2024). FuSIC-KBQA has a three step pipeline: (a) Supervised Retrieval: a supervised retriever, trained on the source domain and optionally fine-tuned on the target domain is used to obtain the top-100 classes, relations and paths that are relevant to the question asked, (c) LLM Generation: We provide the top-10 classes, top-10 relations and top-5 paths along with few-shot exemplars to generate the SPARQL query.
Since no code is available for this model, we use our own implementation based on the description in the paper. For FuSIC-KBQA, and FUn-FuSIC we use LLM $temperature=0$.
A.8.2 Training Details for Supervised Models
We use Hugging Face  Wolf et al. (2020), PyTorch Paszke et al. (2019) for our experiments and use the Freebase setup specified on github  [^8] https://github.com/dki-lab/Freebase-Setup . We use NVIDIA A100 GPU with 40 GB GPU memory and 32 GB RAM. For training the discriminator module of RetinaQA, we require 2 GPUs. (1) For the answerable experiments, we use the supervised models as specified in Patidar et al. (2024). (2) For the unanswerability experiments, we train all models from scratch. (a) We use RnG-KBQA entity linker  [^9] https://github.com/salesforce/rng-kbqa/tree/main/GrailQA/entity_linker (BSD 3-Clause License) trained on the answerable subset of GrailQAbility for all our experiments. (b) We train the RnG-KBQA path retriever on answerable subset of WebQSP [^10] https://github.com/salesforce/rng-kbqa/blob/main/WebQSP/scripts/run_ranker.sh (BSD 3-Clause License). The number of training epochs is determined by the performance of the model over the answerable questions in the dev set. (c) We train the TIARA schema retriever on the answerable subset of WebQSP  [^11] https://github.com/microsoft/KC/tree/main/papers/TIARA/src (MIT License) (d) We train the sketch generator and discriminator of RetinaQA on the answerable subset of WebQSP [^12] https://github.com/dair-iitd/RetinaQA.
A.8.3 Inference Details for Supervised Models
We train all RetinaQA components on the source WebQSP’s training set, using the corresponding target domain’s dev set as a validation set for early stopping. In the absence of unanswerable questions for training, both models use a threshold fine-tuned on a dev set to detect schema-level unanswerability. We again use the target dev sets for this.
We use the dev set in RetinaQA, during discriminator inference for different purposes. (A) Determining how to best utilize the candidate paths. The possibilities are (i) not providing candidate paths, (ii) providing candidate paths in GrailQA format, and (iii) providing candidate paths in WebQSP format. We select the best alternative based upon the performance of the model over the dev set. For the WebQSP $\rightarrow$ GrailQAbility dataset, we observe (ii) works best, whereas for the WebQSP $\rightarrow$ GraphQAbility dataset, we observe (i) works best. (B) Determining the threshold value. RetinaQA applies a threshold on the scores - for a question, if the highest score candidate logical form has a score less than the threshold, the question is labeled as NK. We choose the optimal value of the threshold to maximize the overall EM-s score over the dev set.
A.9 Pangu Adaptation Details
Similar to RetinaQA, we train all Pangu components on WebQSP, using the corresponding target domain’s dev set as a validation set for early stopping. We use one GPU for training. Same as RetinaQA, we use the dev set to determine the threshold for schema-level unanswerability. Pangu-T applies a threshold on the scores - for a question, if the highest score candidate logical form has a score less than the threshold, the question is labeled as NK. We choose the optimal value of the threshold to maximize the overall EM-s score over the dev set.
A.9.1 KB-Binder Adaptation Details
For KB-Binder, we make use of publicly available code  [^13] https://github.com/ltl3A87/KB-BINDER (MIT License). We use self-consistency and majority voting with 6 examples, as in the experiments in the paper. In the retrieval(-R) setting, KB-Binder samples demonstration examples by retrieving from the entire available training data. We restrict its retrieval to our target training set $D_{t}$ with 25 examples. KB-Binder reports experiments using code-davinci-002 as the LLM. For consistency and fair comparison, we replace this with gpt-4-0613 as in other LLM-equipped models. KB-Binder generates logical forms in s-expression, which we preserve.
WebQSP $\rightarrow$ WebQSP $\rightarrow$
Model GrailQAbility GraphQAbility
V3(a) 92 96
V3(b) 94 90
V3(c) 92 94
V3(overall) 88 90
V4(b) 75 68
Table 8: Accuracy of weak verifier on the two datasets by manual analysis of 100 instances from the test sets.
A.10 Accuracy of Weak Verifiers
Accuracy of the two weak verifiers are recorded in detail in Tab. 8.
A.11 FUn-FuSIC Prompts
Here we provide details of various prompts used by FUn-FuSIC.
A.11.1 PUn prompt
The following prompt is for Prompting for Unanswerability (PUn).
Header Prompt Translate the following question to sparql for Freebase based on the candidate sparql, candidate entities, candidate relations and candidate entity types which are separated by "|" respectively. Please do not include any other relations, entities and entity types. Your final sparql can have three scenarios: 1. When you need to just pick from candidate sparql. 2. When you need to extend one of candidate sparql using the candidate relations and entity types. 3. When you will generate a new sparql only using the candidate entities, relations and entity types. For entity type check please use this relation "type.object.type".D o not use entity names in the query. Use specified mids. If it is impossible to construct a query using the provided candidate relations or types, return "NK". Make sure that the original question can be regenerated only using the identified entity types, specific entities and relations.
NK exemplar Question: the tv episode segments spam fall under what subject? Candidate entities: spam m.04vbm Candidate paths: SELECT DISTINCT ?xWHERE ?x0 ns:tv.tv_segment_performance.segment ns:m.04vbm .?x0 ns:tv.tv_segment_performance.segment ?x .?x ns:type.object.type ns:tv.tv_episode_segment . | ... Candidate entity types: tv.tv_series_episode| tv.tv_episode_segment | ... Candidate relations: tv.tv_series_episode.segments (type:tv.tv_series_episode R type:tv.tv_episode_segment)| tv.tv_subject.tv_programs (type:tv.tv_subject R type:tv.tv_program)|... sparql:NK
Question Prompt Question: which school newspaper deals with the same subject as the onion? Candidate entities: the onion m.0hpsvmv Candidate paths: SELECT DISTINCT ?xWHERE ns:m.0hpsvmv ns:book.newspaper.circulation_areas ?x0 .?x0 ns:periodicals.newspapers ?x .?x ns:type.object.type ns:book.newspaper . |... Candidate entity types: education.school_newspaper| type:book.newspaper... Candidate relations: education.school_newspaper.school (type:education.school_newspaper R type:education.educational_institution) | book.newspaper_issue.newspaper (type:book.newspaper_issue R type:book.newspaper)|... sparql:
A.11.2 FUN Prompts
The following is the prompt used by FUn for accommodating feedback from verifier V1.
Syntax error(V1) Feedback Correct the syntax of the following sparql query. Return ONLY the corrected sparql query without any explanation sparql: SELECT ?x AND ?y ... Virtuoso error: word AND not defined
The following is the prompt used by FUn for accommodating feedback from verifier V2.
KB Inconsistency(V2) Feedback The generated sparql has a semantic issue warning: The types of relations don’t match for variable ?x in the query. The assigned relation types by [’computer.computer_emulator.computer’, ’type.object.type computer.computer_peripheral’] are [’computer.computer’, ’computer.computer_peripheral’]. These types are mutually incompatible... Please generate again a different executable sparql using the same context and constraints. DO NOT APOLOGIZE - just return the best you can try.
The following is the prompt used by FUn for accommodating feedback from verifier V3.
Question Logical form disagreement(V3) feedback The question that you answer is NOT same as what you’ve been asked for! You have answered the question "Which opera productions has Gino Marinuzzi conducted?" but you were asked to answer "what is the name of the premiere opera production conducted by gino marinuzzi?". Please generate again a different executable sparql using the relations, classes and entities provided earlier. DO NOT APOLOGIZE - just return the best you can try.
The following three prompts are used by FUn for accommodating feedback from verifier V4.
Answer Inconsistency(V4b) feedback The generated sparql gives an empty answer when executed on freebase KG, Please generate again a different executable sparql using the same context and constraints.
Intermediate Node(V4a) feedback The generated sparql returns an intermediate type node when executed on the freebase KG. Maybe the answer node is an adjacent node to what we currently query for. Please generate again a different executable sparql using the same context and constraints.
Answer Inconsistency(V4a) feedback The logical form upon execution returns International System of Units, which is not answering the question. Please reconstruct the query using same context and constraints.
A.11.3 Prompt for Question Logical Form Agreement Verifier (V3)
The few shots provided for verifying question logical form agreement are derived from $D^{t}$. We obtain positive samples from the dataset $D^{t}$ directly, using the questions and gold logical forms. For obtaining negative samples, we perform zero-shot FuSIC-KBQA inference over $D^{t}$. Then we consider those questions for which the predicted logical form is different from the gold logical form.
First, we perform back-translation to obtain natural language question from the logical form using the following prompt.
Naturalization of variable names(V3(i)) change the sparql query to have variable names representative of what objects they refer to. transform the variable names in this query. Do NOT change the prefix headers and relation names
[…]
We use few-shot LLM prompting to obtain the explanation for question and logical form agreement or disagreement.
Explanation Generation Prompt Explain why the two questions are different. Question we answer: who all like to eat apple or mango? Question originally asked: what are the people who enjoy both apple and mango? explanation: The question we answer returns people. The question originally asked also returns people. The question we answer finds those people who like eating apple, those people who like eating apple. The question originally asked also finds those people who like eating apple, those people who like eating apple. The question we answer uses logical operator OR. However, the question originally asked uses the logical operator AND Hence, they are different. [total 3 exemplars] Question we answer: Which game engines are successors to the Unreal Engine? Question originally asked: which video game engine’s successor engine is unreal engine? explanation:
These few-shots for obtaining the explanation are dataset independent, and are manually written. Some examples of few shots are below.
Question Logical Form Agreement Check(V3(iii)) Question we answer: Who are the cricket players who have made exactly 31 stumps in one day internationals? Question originally asked: name the cricket player who has 31 odi stumps. explanation: The question we answer returns cricket players. The question originally asked also returns cricket players. The question we answer finds cricket players who have made exactly 31 stumps in one day internationals. The question originally asked also finds cricket players who have made 31 stumps in one day internationals. Both questions involve no mathematical or logical operators. Hence, they are same. Question we answer: Which game engines are successors to the Unreal Engine? Question originally asked: which video game engine’s successor engine is unreal engine? explanation: The question we answer returns game engines. The question originally asked also returns game engines. The question we answer finds successors to the Unreal Engine. The question originally asked finds the predecessor of the Unreal Engine. The reasoning steps followed by the two questions are different. Hence, they are different. [total 6 exemplars] Question we answer: Which cars drive at a speed of 80? Question originally asked: name the car with driving speed at least 80? explanation:
A.11.4 scUn Prompt for Logical Form Selection
We store the back-translated natural language questions along with each predicted logical form. We use the following prompt to select the closest back-translated natural language question to the original question.
Best Logical Form Selection Prompt orig_nl_qn = which surf films has sarah finn served as the casting director? 1. pred_nl: Which surfing films has Sarah Finn directed the casting for? 2. pred_nl: Which surfing films has Sarah Finn been the casting director for? of the 2 predicted nl questions, which is closest to the original nl question. Even if none is very close, return the one that is semantically closest? Please explain your answer as well
A.12 Example of FUn Iterations
The following shows 3 Fun iterations for an example question from the test set of WebQSP $\rightarrow$ GrailQAbility dataset.
Question: what is the musical genre of the recording who m i (feat. 일리닛, new champ, myk)?
Entities: who m i (feat. 일리닛, new champ, myk) m.0123lk0s
Logical Form 1 ?x {
    :m.0123lk0s :music.genre.recordings ?x
    ?x :type.object.type :music.genre}
Feedback 1 Type: Logical Form Semantic Error Description: The types of relations don’t match for entity in the query. The assigned relation types by [’music.genre.recordings’] are [’music.genre’]. These types are not associated with this entity in the KB.
Logical Form 2 ?y {
    :m.0123lk0s :music.recording.artist ?x
    ?x :music.artist.genre ?y
    ?y :type.object.type :music.genre}
Feedback 2 Type: Natural Language Semantic Error Description: The question originally asked genre of the song. However, the question we answer returns genre associated with artist of the song.
Logical Form 3 ?x { ?x
    :music.genre.recordings :m.0123lk0s
    ?x :type.object.type :music.genre}
Feedback 3 Status: Passes All Checks Details: lf3 satisfies syntax, logical form semantic, natural language semantic and execution checks.